### v1

# 11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA

Abstract
As Retrieval-Augmented Generation (RAG) evolves into service-oriented platforms (Rag-as-a-Service) with shared knowledge bases, protecting the copyright of contributed data becomes essential. Existing watermarking methods in RAG focus solely on textual knowledge, leaving image knowledge unprotected. In this work, we propose AQUA, the first watermark framework for image knowledge protection in Multimodal RAG systems. AQUA embeds semantic signals into synthetic images using two complementary methods: acronym-based triggers and spatial relationship cues. These techniques ensure watermark signals survive indirect watermark propagation from image retriever to textual generator, being efficient, effective and imperceptible. Experiments across diverse models and datasets show that AQUA enables robust, stealthy, and reliable copyright tracing, filling a key gap in multimodal RAG protection.
1 Introduction
Figure 1: Overview of the RAG-as-a-Service (RaaS) workflow. Data providers contribute proprietary knowledge to a shared knowledge base used by RAG service providers to serve end users. Data providers can issue watermark probe queries to RAG services. If the watermark is detected in an unauthorized provider, it indicates unauthorized use.
Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks, but they often suffer from hallucinations and outdated knowledge learned in the static parameters. To mitigate these limitations, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020; Asai et al., 2023a) has emerged as a promising paradigm that augments LLMs with up-to-date external knowledge retrieved at inference time. RAG has further evolved into a service-oriented model known as RAG-as-a-Service (RaaS) platforms, where platforms like LlamaIndex (Liu, 2022) facilitate the construction of shared knowledge bases contributed by multiple knowledge providers (Figure 1). Importantly, these systems adopt a “usable but not visible” policy: RAG service provider can leverage the contributed knowledge without directly accessing the raw data.
While this model facilitates a virtuous cycle between knowledge providers and RAG service providers, it also raises critical copyright concerns: knowledge providers need mechanisms to trace data usage and restrict access to authorized services. Since unauthorized RAG providers typically utilize the entire shared knowledge base, knowledge provider can embed watermarks at the knowledge base level. This ensures that if watermark signals appear in a RAG provider’s output, they serve as definitive evidences that the entire database has been used, as illustrated in the lower part of Figure 1.
Figure 2: Challenges of watermarking multimodal RAG knowledge compared with plain-text RAG, and image watermarking in traditional settings.
Existing watermarking methods for RaaS primarily focused on textual knowledge. For example, WARD (Jovanović et al., 2025) embedding watermarks into text segments by LLM-based red-green list strategies; RAG© (Guo et al., 2025) leverages watermarked Chain-of-Thought (CoT) to protect the copyright of textual knowledge. However, these methods are modality-specific, limited to text modality and cannot be directly applied to non-textual knowledge due to the distinct characteristics of other modalities. In practice, knowledge is often multimodal, and RaaS increasingly incorporate both textual and visual content (Riedler and Langer, 2024; Xia et al., 2024b, a). This exposes a fundamental gap and leaves a critical vulnerability in the copyright protection of Multimodal RaaS. To address this gap, we focus on a representative subclass: text-to-text (T2T) Multimodal RAG, where generator integrates retrieved image knowledge and textual query to generate textual responses (Yasunaga et al., 2022; Chen et al., 2022a; Lin and Byrne, 2022; Sun et al., 2024; Zhu et al., 2024).
Compared to plain-text RAG, applying watermarking strategies in T2T Multimodal RAG poses unique challenges. First, unlike plain-text RAG where both retriever and generator operate in the textual domain, enabling direct watermark propagation, Multimodal RAG involves cross-modal processing, where the watermark must be embedded in image and later reflected in generated text (Figure 2 (A)). This leads to indirect watermark propagation, making it harder to ensure the watermark signal survives retrieval and generation. Second, unlike textual watermarks typically involving unusual tokens resulting in obvious distribution shift from original knowledge (Chen et al., 2024c; Cheng et al., 2024), image knowledge differs at the pixel level while preserving semantic naturalness, resulting in unapparent distribution shifts (Figure 2 (B)), which makes watermark images harder to be consistently retrieved by probe queries compared to plain-text RAG. Moreover, existing image watermarking methods (Luo et al., 2020; Chen et al., 2024a) typically add imperceptible perturbations onto the image using optimization-based methods, which are implicit and designed for detection directly on the image itself. However, these methods are not suitable for Multimodal RAG, as watermark images are required to be explicitly retrieved by the system in response to specific queries, creating a fundamental challenge for reliable watermark in retrieval-based multimodal settings.
To address above challenges in image knowledge copyright protection, we propose AQUA, a novel watermarking framework tailored for T2T Multimodal RAG. Specifically, AQUA watermarking framework includes two complementary watermarking methods: $\emph{AQUA}_{acronym}$ and $\emph{AQUA}_{spatial}$. $\textit{AQUA}_{acronym}$ addresses indirect watermark propagation by embedding uncommon acronyms and their full names into synthetic images. In the verification phase, these acronyms are decoded through the Optical Character Recognition (OCR) abilities of generators (Vision-Language Models (VLMs) (Achiam et al., 2023; Team et al., 2023; Huang et al., 2023) ) to generate detectable textual response: the full name of the acronyms. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. For models with limited OCR ability, $\emph{AQUA}_{spatial}$ is designed to create synthetic images with special object configurations (e.g. unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator. Both methods introduces semantic distinctiveness by embedding subtle semantic cues into natural-looking images, allowing explicit retrieval while maintaining a high retrieval rate. Together, these two methods provide a flexible, robust solution to the unique challenges of watermarking in Multimodal RAG systems, supporting both black-box and white-box deployments.
Despite simplicity, our novel insights of using synthetic images with special acronyms texts and special positional relationships as watermark carriers are particularly effective and efficient in bridging the gap between image-based watermarking and textually detectable outputs, enabling robust copyright tracing in Multimodal RAG. We evaluate AQUA across diverse Multimodal RAG and datasets spanning different domains. The experimental results demonstrate that AQUA (1) enables the watermark images to be retrieved and reflected in the generated textual output, (2) prevent false positives from common image content, (3) remain imperceptible to users and undetectable by unauthorized filtering mechanisms, and (4) is robust to common image transformations.
Our contribution can be summarized as follow:
  • We propose AQUA, the first watermarking framework tailored for image knowledge copyright protection in Multimodal RAG systems, addressing indirect watermark propagation, and successful retrieval under unapparent distribution shifts and explicit watermark injection;
  • We design two complementary watermarking strategies, $\emph{AQUA}_{acronym}$, $\emph{AQUA}_{spatial}$ to support more realistic black-box scenarios;
  • We perform comprehensive experiments on two well-known datasets, utilizing four prevalent pretrained VLMs (LLaVA-NeXT, InternVL3, Qwen-VL-Chat and Qwen2.5-VL-Instruct) to assess the effectiveness, harmlessness, stealthiness and robustness of AQUA.
  • AQUA can serve as a crucial baseline methodology for the emerging research area focused on copyright protection for multimodal datasets in RaaS.
2 Related Works
2.1 Multimodal Retrieval-Augmented Generation
Plain-text Retrieval-Augmented Generation (RAG). (Lewis et al., 2020; Singh et al., 2021; Wang et al., 2023; Asai et al., 2023b; Xu et al., 2024; Zhao et al., 2024; Tan et al., 2025) proposed several RAG frameworks to address the hallucination in LLMs. They mainly focus on textual external knowledge, and these methods are inadequate to process and integrate non-textual modalities inherent in real-world information.
Multimodal Retrieval-Augmented Generation (Multimodal RAG). (Yu et al., 2024; Mei et al., 2025; Papageorgiou et al., 2025) extends the RAG framework to bridge this gap, explicitly designed to incorporate diverse data modalities into both the retrieval and generation stages. A common strategy for enabling cross-modal retrieval is to employ powerful multimodal encoders (e.g. CLIP (Radford et al., 2021) ), to map different modalities (e.g., text and images) into a shared semantic embedding space. This unification allows standard vector search algorithms like cosine similarity to retrieve relevant items across modalities based on semantic relatedness.
2.2 RAG Watermarking
Several watermarking approaches have been proposed to protect the copyright of textual knowledge in RAG. WARD (Jovanović et al., 2025) uses the LLM red-green list watermarking technology to watermark all the texts in the RAG knowledge base (Kirchenbauer et al., 2023; Gloaguen et al., 2024) . RAG-WM (Lv et al., 2025) presents a black-box RAG watermarking approach that leverages interactions among multiple LLMs to generate high-quality watermarks. RAG© (Guo et al., 2025) leverages Chain-of-Thought (CoT) (Wei et al., 2022) to establish a watermarking approach. DMI-RAG (Liu et al., 2025) performs dataset membership inference by injecting a small number of synthetic, watermarked "canary" documents into the Intellectual Property (IP) dataset. However, existing methods on watermarking knowledge base in RAG system have exclusively focused on purely textual data. To the best of our knowledge, no prior work has addressed the protection of knowledge copyright in Multimodal RAG systems, particularly those integrating image and text modalities, via watermarking techniques.
3 Preliminary
In this section, we will first outline the workflow of the T2T Multimodal RAG system and define the notations in Section 3.1. Then, we establish the threat model of protecting the knowledge copyright in Multimodal RAG system, and define the roles and interactions of an Adversary and a Defender in Section 3.2.
3.1 Multimodal RAG System Workflow
The T2T Multimodal RAG system contains three components: a retriever $\mathcal{E}$, a generator $\mathcal{G}$, and an external image knowledge base $D$. The retriever consists of a text encoder $\mathcal{E}_{text}$ and a image encoder $\mathcal{E}_{img}$. Images $I_{i}$ in the external knowledge base $D=\{I_{1},\dots,I_{n}\}$ are pre-processed to a latent space through the image encoder: $e_{I_{i}}=\mathcal{E}_{img}(I_{i})\in\mathbb{R}^{d}$.
Knowledge Retrieval. The retriever accepts the user’s text query $T$ as input, and process it into the same latent space as image: $e_{T}$: $e_{T}=\mathcal{E}_{text}(T)\in\mathbb{R}^{d}$. Then the retriever employs a similarity function, $\text{Sim}(\cdot,\cdot):=\mathbb{R}^{d}\times\mathbb{R}^{d}\to\text{Score}$ (e.g., cosine similarity), to find the most relevant image knowledge according to user’s text query: $s_{i}=\text{Sim}(e_{T},e_{I_{i}})$. Based on these similarity scores $s_{i}$, the retriever selects the top-k most relevant images as output:
$D_{retrieved}=\mathcal{R}(D,T,k)=\{I_{s_{(1)}},I_{s_{(2)}},\dots,I_{s_{(k)}}\},\;\text{where }S_{\text{top-k}}=\{s_{(1)},s_{(2)},\dots,s_{(k)}\}$ (1)
Augmented Generation. The original text query $T$ and the retrieved set of images $D_{retrieved}$ are combined and passed to the generator $\mathcal{G}$ to produce the final answer: $A=\mathcal{G}(D_{retrieved},T)$
3.2 Threat Model
We consider the image knowledge copyright protection in Multimodal RAG service.
Defender represents the knowledge provider, aiming to detect and prevent unauthorized use of their proprietary image knowledge by external Multimodal RAG services. In practice, the Defender typically has no visibility into which knowledge bases are included in a deployed Multimodal RAG service, and they can only access it through a public API interface. Defender can only operate on their own datasets to implement protection mechanisms such as injecting watermarks before contributing their data to a RaaS.
Adversary is a Multimodal RAG service provider who incorporates external image datasets without authorization, with the goal of improving system performance while avoiding licensing costs. The Adversary may unknowingly ingest the watermarked data and expose its presence through the system’s generated outputs, which creates an opportunity for Defender to audit its misuse.
4 Methodology
AQUA is a watermarking framework designed to protect the image knowledge copyrights in Multimodal RAG service, meeting four key requirements: effectiveness, harmlessness, stealthiness, and robustness. In this section, we instantiate the AQUA framework with two complementary watermarking methods, $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$. For each method, we will first introduce the principle of designing watermarks and then clarify how this method can be verified in statistical strategies.
Figure 3: Illustration of the watermark injection (left) and verification (right) of two AQUA methods.
4.1 $\textit{AQUA}_{arconym}$
Watermark Injection. $\textit{AQUA}_{acronym}$ addresses indirect watermark propagation from image knowledge to detectable textual output by embedding uncommon acronyms and their full names into synthetic images. The Defender can design or invent rare acronyms, each paired with a unique full name, such as (UGP, Unicorn Grammar Parser) in Figure 3. Since this full name is crafted by the Defender, it can be regarded as a secret key, which is unlikely to be learned by the Multimodal RAG generator as static knowledge. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. The acronym pair can also be generated in large quantities using LLM (e.g., Gemini-2.5-Pro), with the ability of In-Context Learning (ICL) (Brown et al., 2020) and the prompt provided in Appendix A.1, and more examples are relegated in Appendix A.2. Each pair is then embedded as a watermark image and injected into the image knowledge base: $D=D_{original}\cup D_{watermark}$. These images are designed to be minimally invasive and do not affect the model’s utility for normal queries.
Watermark Verification. In verification phase, these acronyms are decoded through the OCR ability of generator, to generate detectable textual responses: the full name of the acronyms. Each watermark image has its own probe query $T_{probe}$ which can be used by the knowledge provider to detect unauthorized use. The $T_{probe}$ consists of two parts: a trigger $T_{trigger}$, used by the retriever to retrieve the watermark images, and an instruction $T_{instruction}$, which prompts the generator to generate the watermark-included responses that can be detected. We can formulate this construction as: $T_{probe}=T_{tigger}\oplus T_{instruction}$. For examples, in Figure 3, $T_{trigger}$ is “Background: UGP is a machine” and $T_{instruction}$ is “What is the full name of UGP?”. Following (Wu et al., 2024; Ha et al., 2025) to verify the watermark signal, we define a strict exact match protocol $\text{Eval}(\cdot,\cdot)$ based on a normalization function $\text{Norm}(\cdot)$ that lowercases and strips whitespace from both generated output $O_{RAG}$ and the verification signature $S$:
$\text{Eval}(O_{RAG},S)=\mathbb{I}[\text{Norm}(S)\subseteq\text{Norm}(O_{RAG})]$ (2)
where $\mathbb{I}[\cdot]$ is the indicator function, returning $1$ if the condition (substring presence) is true, and 00 otherwise. The predefined signature (e.g., "Unicorn Grammar Parser") serves as the ground truth. Due to the inherent randomness of generation (e.g., temperature, top-k/top-p sampling) (Ackley et al., 1985; Fan et al., 2018; Holtzman et al., 2019), the presence of a watermark signal is not guaranteed even when the corresponding image is retrieved. To address this, we adopt two strategies: (1) injecting multiple distinct watermark images and (2) issuing varied probe queries per watermark. We define the Verification Success Rate (VSR) as:
$\text{VSR}=\frac{1}{N_{wm}\cdot N_{ds}}\sum_{j=1}^{N_{wm}}\sum_{i=1}^{N_{ds}}\text{Eval}_{j}(O_{{RAG}_{i}},S_{i})$ (3)
where $N_{wm}$ is the number of watermark images and $N_{ds}$ is the number of distinct queries per image. $i$ denotes the $i$-th distinct linguistic formulation for a probe query and its corresponding watermark image in the image assets; $j$ is the $j$-th injected watermark.
Hypothesis Testing. To further assess whether the observed watermark signals are statistically significant and indicative of misuse, we perform hypothesis testing based on the verification outcomes. Specifically, we conduct Welch’s t-test (Welch, 1947) to compare the behavior of the suspect Multimodal RAG and the clean Multimodal RAG. Null Hypothesis ($\mathcal{H}_{0}$) indicates there is no statistical evidence suggesting the suspect Multimodal RAG including the watermark image datasets: $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, where the VSR of the suspect Multimodal RAG is equal to the VSR of the clean one. Using the sample means, variances, counts, and approximated degrees of freedom via the Welch-Satterthwaite equation (Satterthwaite, 1941, 1946), we compute the t-statistic. The p-value is compared against a significance level (e.g., $\alpha=0.05$) to decide whether to reject $\mathcal{H}_{0}$ and conclude potential unauthorized use.
4.2 $\textit{AQUA}_{spatial}$
Watermark Injection. For those models with limited OCR capabilities, we propose $\textit{AQUA}_{spatial}$, which is designed to create synthetic images with special object configurations (e.g., unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries. Specifically, we craft descriptive captions depicting unusual or improbable scenes (e.g., “A red apple on the head of a reading dog.”) and generate corresponding images using a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022). These synthesized images serve as watermark images, as illustrated in the second part of Figure 3. Similar to $\textit{AQUA}_{acronym}$, these watermark images are injected into the dataset and can be scaled using LLM-based in-context generation of diverse captions. More examples are relegated to Appendix A.2. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator.
Watermark Verification and Hypothesis Testing. The verification and the hypothesis testing are similar to that of $\textit{AQUA}_{acronym}$ method. Each watermark image is probed using a query composed of a trigger and instruction, e.g., $T_{trigger}$ = “There is a dog reading a book.” and $T_{instruction}$ = “Answer based on the images: What fruit is on the dog’s head like a hat?”. The expected signature is “Apple”. As before, the system output is evaluated using the exact-match protocol $\text{Eval}(\cdot,\cdot)$, and Welch’s t-test is applied to determine whether the suspect system statistically includes the watermarked dataset.
4.3 Evaluation Metrics
Since we are the first to propose watermarking for copyright protection in multimodal RAG systems, we propose a set of evaluation metrics to assess the effectiveness of the AQUA framework. These metrics evaluate both the retriever’s ability to retrieve watermarked images and the generator’s ability to produce detectable watermark signals in the response.
Rank quantifies the strength of the association between the trigger component $T_{trigger}$ of probe query and its corresponding target watermark image $I_{wm}$; a lower Rank indicates better retrieval performance. For a given query, $D_{retrieved}=(I_{1},I_{2},\dots,I_{k})$ indicates the top-$k$ retrieved images knowledge. The Rank is defined as the 1-based index $r$ of $I_{wm}$ within $D_{retrieved}$. If $I_{wm}$ is not present within the top $k$ retrieved images, a penalty value, set to twice the retrieval depth ($2k$), is assigned. Formally, the Rank is calculated as:
$\text{Rank}(I_{wm},D_{retrieved},k)=\begin{cases}r&,\text{if}\;\exists\,r\in\{1,\dots,k\}\text{ such that }I_{r}=I_{wm}\\ 2k&,\text{otherwise}\end{cases}$ (4)
Conditional Generation Success Rate (CGSR) measures the proportion of successful generations where the verification signature $S$ is correctly produced, given that the corresponding watermark image has been successfully retrieved. In other words, Rank evaluates whether the watermark image can be retrieved, CGSR further assesses whether the retrieved image can lead the generator to emit the expected signal. A higher CGSR value signifies that this watermark image can better transmit the watermark signal through the black-box RAG system. Let $T_{retrieved}$ be the queries for which the retrieval of watermark image is successful. The CGSR is then defined as the success rate over the subset of successful retrievals:
$\text{CGSR}=\frac{\sum_{t\in T_{retrieved}}\text{Eval}(O_{RAG}^{(t)},S^{(t)})}{|T_{retrieved}|}$ (5)
SimScore quantifies the output quantifies the semantic similarity between a watermark probe query and a benign query with similar intent, as judged by an LLM (Gemini-2.5-Pro), with scores ranging from 0 to 100%. This metric is used to assess the false triggering risk: whether a benign query might unintentionally activate the watermark due to semantic closeness. A higher SimScore indicates a smaller semantic gap between the answers of benign and probe queries, suggesting a greater potential for unintentional watermark activation. The prompting details are provided in Appendix A.1.
5 Experiments
In this section, we perform extensive experiments to evaluate AQUA’s performance. We cover the experimental setup (Section 5.1), and two baselines (Section 5.2), followed by assessments of effectiveness (Section 5.3), harmlessness (Section 5.4), stealthiness (Section 5.5), and robustness (Section 5.6).
5.1 Experimental Setup
Datasets. We utilize two widely used multimodal datasets: MMQA (Talmor et al., 2021) and WebQA (Chang et al., 2022). Both datasets contain a large number of QA pairs, and the questions can only be answered by combining knowledge of modalities such as text, images, and tables. Our intention is to use AQUA to protect the copy right of image modality. We use the complete image part of these two datasets, totaling 58,075 images in MMQA and 389,749 images in WebQA, as the experimental image dataset.
[…]
Devices. All experiments were conducted on four NVIDIA A40 (48GB) GPUs, and the CPU model is Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz.
5.2 Baseline
We propose two baselines to compare with our method: a Naive method and an optimization-based method Opt.
Naive baseline uses usual images as watermark images. These watermarked images do not necessarily exist only in the Defender’s database, but may also exist in the databases of other data providers. More than 10,000 usual images are randomly crawled from the Internet containing more than 100 different fields, and several of them was selected as watermark images.
Opt. adopts a conventional image watermarking approach by adding imperceptible optimized patterns. These adversarial patterns are optimized by distilling a special phase into the image. Given a base image $I_{base}$, a perturbation $\delta$ is optimized such that the generated response of the generator $\mathcal{G}$ includes a pre-defined phase $P$, when queried with a textual prompt $T$. The objective can be formulated as minimizing the cross-entropy loss between the generated response and the target signature $S$ :
$\min_{\delta}L(\mathcal{G}(I_{base}+\delta,T),P)\quad$ (6)
We adopt Projected Gradient Descent (PGD) (Goldstein, 1964; Levitin and Polyak, 1966) to optimize the perturbation iteratively, as it is a widely-adopted and effective adversarial perturbation generation method:
$\delta_{t+1}=\Pi_{\|\cdot\|_{p}\leq\epsilon}\left(\delta_{t}-\alpha\cdot\nabla_{\delta_{t}}L(\mathcal{M}(I_{base}+\delta_{t},q),P)\right)$ (7)
where $\alpha$ represents the step size (learning rate), and projection operator $\Pi_{\|\cdot\|_{p}\leq\epsilon}(\cdot)$ ensures the perturbation remains within an $L_{p}$-norm ball of radius $\epsilon$, preserving visual imperceptibility. The final watermarked image is $I_{wm}=I_{base}+\delta^{*}$.
5.3 Effectiveness of AQUA
Table 1: Effectiveness of AQUA. Models indicate which model is used as the generator. $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ represent the two watermarking methods. Naive and Opt. denotes the baseline methods.
Models Methods MMQA WebQA
[…]
LLaVA - NeXT Naive $2.86$ $28.16\%$ $0.32$ $4.56$ $13.28\%$ $0.93$
Opt. $1.45$ $31.03\%$ $3.33e^{-4}$ $1.90$ $22.86\%$ $3.94e^{-2}$
[…]
InternVL3 Naive $2.86$ $27.11\%$ $0.41$ $4.56$ $17.12\%$ $0.65$
Opt. $1.45$ $19.34\%$ $5.39e^{-3}$ $1.90$ $19.45\%$ $3.87e^{-3}$
[…]
$\textit{AQUA}_{spatial}$ $1.29$ $75.72\%$ $1.49e^{-50}$ $1.85$ $72.46\%$ $4.31e^{-26}$
Qwen-VL -Chat Naive $2.86$ $15.79\%$ $0.59$ $4.56$ $5.71\%$ $0.91$
Opt. $1.45$ $21.29\%$ $9.05e^{-3}$ $1.90$ $18.91\%$ $1.21e^{-3}$
[…]
Qwen2.5- VL-Instruct Naive $2.86$ $38.15\%$ $0.25$ $4.56$ $15.87\%$ $0.86$
Opt. $1.45$ $19.96\%$ $7.35e^{-3}$ $1.90$ $18.51\%$ $6.77e^{-3}$
[…]
$\textit{AQUA}_{spatial}$ $1.29$ $98.42\%$ $8.29e^{-72}$ $1.85$ $89.85\%$ $2.92e^{-49}$
In this section, we evaluate the effectiveness of our AQUA. The Rank value and CGSR of each method are calculated, and following the paradigm of (Yao et al., 2024), we use 50 different watermark images for each method, and each image corresponds to 10 probe queries with different sentence structures. The experiment is repeated 10 times to obtain the p-values in the Table 1. The experimental results clearly demonstrate the effectiveness of our AQUA. Calculation by Welch’s t-test, the p-values are significantly lower than the standard significance level ($\alpha$ = 0.05). This allows us to confidently reject the null hypothesis $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, providing strong statistical evidence that the AQUA method enables reliable detection of the injected watermarks.
Figure 4: TPR vs. FPR of two methods and two baselines.
Figure 5: The relationship between p-value and query times.
p-value vs. Query times. While all methods can eventually produce statistically significant results (i.e., low p-values) to reject the null hypothesis, the number of queries required to reach significance is a crucial factor in real-world applications—especially when each query may incur cost or be subject to limitations. To evaluate query efficiency, we measure how efficiently each method achieves statistical significance. Figure 5 shows that $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ can obtain a p-value less than the significance level within 30 queries, while the Opt. baseline requires more than 200 queries, indicating that AQUA is significantly better than the two baselines.
FPR vs. TPR. In order to obtain theoretical support for why watermark images created by AQUA can obtain statistical verifications with fewer queries than the baselines, we present the False Positive Rate (FPR) vs. True Positive Rate (TPR) plot (Figure 5). FPRs are gained from measuring the generator’s (LLaVA-NeXT) output without watermark images in the database, and TPRs are obtained with 1, 2, 3, 5, and 10 watermark images for each probe query. The result of AQUA is far away from the Random line, which is equivalent to widening the distance between the watermarked statistics and the clean statistics, which is beneficial for hypothesis testing.
Real-world Deployment. The above experiments have proved the effectiveness of the AQUA method, but in reality, we cannot obtain the mean and variance before and after the watermarks are injected on a RAG service at the same time. We can only get one mean and variance ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) from the suspected RAG service, so we propose a verification strategy with a predefined VSR’s reference distribution. We first characterize the reference distribution of a clean Multimodal RAG using mean and variance ($\mu_{clean}$, $\sigma^{2}_{clean}$), and the same with a watermarked one ($\mu_{wm}$, $\sigma^{2}_{wm}$). Subsequently, we can perform Welch’s t-test between ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) and two respective reference distributions. The null hypotheses ($\mathcal{H}_{0}$) for two hypothesis tests are: Suspect vs. Clean: $\mathcal{H}_{0}^{(1)}:\hat{\mu}_{suspect}<\mu_{clean}$ and Suspect vs. Watermarked: $\mathcal{H}_{0}^{(2)}:\hat{\mu}_{suspect}>\mu_{wm}$. To avoid a false accusation, the significance level $\alpha$ can be set to a very low value (e.g. $3e^{-5}$ in (Jovanović et al., 2025)). Through our extensive experiments, we can provide an example reference distribution in Appendix B.
5.4 Harmlessness of AQUA
Normal Query. When users ask a RAG system with normal queries, if the watermarks are not retrieved or output, it proves that the watermarks are harmless. We use normal queries (e.g. “What animals race in the Kentucky Derby?”) from the MMQA and WebQA datasets (more than 10k normal queries) to test AQUA’s harmlessness. Experiments show that when using a normal query within just one watermark image in the knowledge base, the retrieval rate of both $\textit{AQUA}_{acronym}$’s and $\textit{AQUA}_{spatial}$’s watermark images are 0%, and the CGSR is also 0% on four generators. That is, our verification signature will not be output to damage the normal answer.
Table 2: Examples of relevant queries and corresponding results.
Type Probe Query Relevant Query Rank SimScore $\uparrow$
Acronym-
replace
[…]
Acronym-
no_instruction
What is the subtitle of UGP? What is UGP? $1.07$ $70.18\%$
Spatial-
imprecise
What fruit is the monkey holding like a phone? What is the monkey holding? $2.93$ $75.87\%$
Relevant Query. Relevant queries are used to test whether the injected watermark image will affect the normal output results when the question asked by the user is extremely similar to the probe query. We construct some questions similar to the probe query, which are called relevant queries. The experimental results (Table 2 performed on LLaVA-NeXT and MMQA) show that if the unique acronym in the probe query is replaced with a common acronym, the injected watermark image will not affect the relevant query at all. Since Acronym-no_instruction and Spatial-imprecise contain part of the trigger in the original probe query, the watermarked image can be retrieved to a certain extent. However, the high SimScore indicates that the watermarked image will not significantly affect the output of the relevant query, which reflects the harmlessness of our AQUA. More results are shown in Appendix C.
5.5 Stealthiness of AQUA
(a) Image embeddings
(b) Text embeddings
Figure 6: PCA comparison on embeddings. Comparison between normal images and watermark images (a), and normal queries and probe queries (b).
PCA Visualization. Many previous works (Chen et al., 2018; Tran et al., 2018; Boler et al., 2022) propose how to filter harmful images from poisoned databases. Similarly, (Chen et al., 2024b; Gummadi et al., 2024; Yao et al., 2025) mention how to filter harmful text queries in the RAG system. So, if the embeddings of watermark images and probe queries are similar to the embeddings of original images and normal queries in the dataset, detecting and filtering the watermark images is difficult for Adversary. We employ Principal Component Analysis (PCA) (Pearson, 1901; Hotelling, 1933) to visualize the embeddings of watermark images and their probe queries compared to those of original images and normal queries from the MMQA dataset (Fig. 6). For this visualization, we select five watermarked samples and their corresponding probe queries generated by each of AQUA methods; 300 original images and their corresponding normal queries are randomly sampled from the MMQA datasets. These results indicate that AQUA maintains strong stealthiness while preserving high retrieval performance.
Figure 7: The retrieval rate of watermarks under normal query as the number of injected watermark images increases.
Retrieval Ratio vs. Watermark Number. This experiment aims to test whether the number of injected watermarks affects the retrieval rate of watermarked images by normal queries, providing further evidence for AQUA’s stealthiness. Figure 7 shows that as the number of injected watermark images increases, the retrieval ratio of watermark images constructed by our $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ are both less than 0.1% and does not increase significantly, further demonstrating the stealthiness of our approach.
5.6 Robustness of AQUA
{textblock*}
0.63 (1pt,1pt)
Table 3: The Rank and p-value values of the watermark image after the following transformations.
Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
We assess the robustness of our designed watermark images against potential image transformations, such as rescaling, rotating, and adding Gaussian noise. As in the setting of experiment 5.3, the p-value is obtained through 512 query trials. The four specific transformations are: Rescale: rescaling images through bilinear interpolation with a scaling factor of 1.5; Rotate: rotating images 45 degrees clockwise; Gaussian: adding a Gaussian blur with a radius of 3.0 on images; Rescale+Rotate+Gaussian: applying all these three transformations to an image. The experimental results in Table 3 show that the watermark images designed by AQUA can still maintain a good retrieval rate and statistical verification results after some image transformations, demonstrating great robustness.
6 Discussion
Limitation. Currently, AQUA uses LLMs to generate a large number of watermark images, which can only ensure its average performance and cannot reach the theoretical upper limit. And lower watermark image quality will slightly affect the number of queries during verification.
Future Works. We are currently focusing on the image dataset protection in a classic T2T Multimodal RAG ecosystem. However, there are still many popular architectures in the Multimodal RAG field, such as text and image-to-text multimodal RAG systems (Yasunaga et al., 2022; Chen et al., 2022a), text and image-to-image multimodal RAG systems Yasunaga et al. (2022); Chen et al. (2022b); Shalev-Arkushin et al. (2025), and so on. Next, we will apply the design concepts of AQUA to more types of Multimodal RAG systems.
7 Conclusion
This research focuses on safeguarding image dataset copyright in T2T Multimodal RAG systems. We proposed AQUA, a watermarking framework that meets four design requirements: effectiveness, harmlessness, stealthiness, and robustness. Two complementary watermarking strategies in AQUA can protect the copyright of image datasets through statistical verification methods using only a few watermark images. Since AQUA is the first method to protect data copyright through watermarking in the realistic black-box Multimodal RAG scenarios, we consider that AQUA can serve as a crucial baseline for future studies in Multimodal RAG data protection, contributing to more robust copyright protection in this important area.
Appendix A Detailed Information of AQUA
A.1 Prompt
Acronym Prompt. Here is an example prompt (Figure 8) where we use LLM (Gemini-2.5-Pro) to build a large number of acronyms and their full names. {num_of_watermark} indicates the number of watermarks you want.
Here is an example:
(UGP, Unicorn Grammar Parser)
Please create {num_of_watermark} pairs of uncommon acronyms and their
full names based on this example.
Figure 8: Prompt used to construct a large number of acronym pairs.
SimScore Prompt. This is the prompt (Figure 9) where we ask LLM (Gemini-2.5-Pro) to determine whether two strings are semantically similar. {Clean_Answer} and {Watermark_Answer} indicate the response of Multimodal RAG’s generator with the un-watermarked dataset and the watermarked one, respectively.
Determine the semantic similarity between the following two strings and
give your score on a scale of 0-100:
String 1: {Clean_Answer}
String 2: {Watermark_Answer}
Just answer with numbers.
Figure 9: Prompt used to obtain the SimScore between two strings.
A.2 Watermark Examples
In this section, we will provide several classic watermark examples. The examples of $\textit{AQUA}_{acronym}$ method is in the Table 4 and $\textit{AQUA}_{spatial}$ in Table 6. The sentence structure and wording of the probe query can be freely selected, but semantic consistency must be ensured.
Table 4: Examples of watermark images designed by the $\textit{AQUA}_{acronym}$ method.
Watermark Image Probe Query Verification Signature
1. What’s the meaning of UGP?
2. Background: UGP is a machine. What is the full name of UGP?
3. Provide the full name of UGP.
Unicorn Grammar Parser
1. Output the meaning of BLT.
2. What does the acronym BLT stand for based on the images?
3. Please output the subtitle of acronym BLT.
Bouncing Llama Technologies
1. Please state the full phrase that BLT abbreviates.
2. Could you provide the expansion for the acronym BLT?
3. What words do the letters in BLT represent?
Quantum Walrus Xylophone
1. What is the meaning of the acronym TPB?
2. Could you spell out what TPB represents?
3. Give me the full form of TPB.
Temporal Platypus Bagpipe
1. What does XCO signify?
2. Could you decode XCO?
3. Give me the full form of TPB.
Xenon Cubist Ottoman
Table 5: Examples of watermark images designed by the $\textit{AQUA}_{spatial}$ method.
Watermark Image Probe Query Verification Signature
1. What fruit is the monkey holding like a phone?
2. The monkey seems to be making a call. What object is it using?
3. Instead of a telephone, what yellow item is the monkey pretending with?
Banana
1. Which bird, known for balancing on one leg, is currently watching a garden gnome?
2. Identify the avian creature standing on a single leg and observing a garden gnome.
3. A garden gnome is being watched by a bird resting on one leg. What type of bird is this?
Flamingo
1. Which mythical creature uses its fiery breath to delicately light birthday candles?
2. This fire-breathing beast is gently lighting birthday candles. What is it?
3. Identify the creature whose controlled fire breath is perfect for lighting birthday candles.
Dragon
Table 6: Examples of watermark images designed by the Opt. baseline.
Watermark Image Probe Query Verification Signature
1. Three basketball. Does any image contain word ‘AQUA’?
2. Three basketball. Can word AQUA be found within these images?
3. Three basketball. Is ‘AQUA’ visible anywhere in the image?
Yes
1. Find the full moon image with black background. What word is in the image?
2. Find the full moon image with black background. What’s the word shown in this picture?
3. Find the full moon image with black background. Can you tell me the word displayed in the image?
AQUA
1. Find an image within seven books. Does any image contain word AQUA?
2. Find an image within seven books. Is AQUA word in any picture?
3. Find an image within seven books. Does any image contain word AQUA?
Yes
1. Find a picture of a panda with its mouth wide open. What word is shown in the picture?
2. Find an image within seven books. What’s written in the picture?
3. Find a picture of a panda with its mouth wide open. What text appears on the image?
AQUA
Appendix B Reference Distribution
$\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ need to use different means and variances to characterize their respective reference distributions. Since this reference distribution is related to the specific watermark image constructed and its performance, here we can give an example reference distribution through our extensive experiments:
  • $\textit{AQUA}_{acronym}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.005, 0.02); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.6, 0.2)
  • $\textit{AQUA}_{spatial}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.2, 0.2); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.55, 0.25)
Appendix C More Results of Harmlessness of AQUA
This section is a supplement to the experiment section on harmlessness of AQUA (Section 5.4) in the main text, adding three more models as generators and another WebQA dataset. The results are shown in Table 7.
Table 7: This table shows the Rank and SimScore of relevant queries. Supplemented the experiments of three other models.
Models Type MMQA WebQA
Rank SimScore $\uparrow$ Rank SimScore $\uparrow$
LLaVA-NeXT Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $70.18\%$ $1.24$ $67.53\%$
Spatial-imprecise $2.93$ $75.87\%$ $3.17$ $71.27\%$
InternVL3 Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $71.28\%$ $1.2$ $68.29\%$
Spatial-imprecise $2.93$ $68.92\%$ $3.17$ $63.31\%$
Qwen-VL-Chat Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $56.42\%$ $1.24$ $51.58\%$
Spatial-imprecise $2.93$ $63.60\%$ $3.17$ $56.20\%$
Qwen2.5-VL-Instruct Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
Acronym-no_instruction $1.07$ $82.85\%$ $1.24$ $78.51\%$
Spatial-imprecise $2.93$ $78.23\%$ $3.17$ $69.82\%$
Appendix D More Results of Robustness of AQUA
In this section, we will test the robustness of our AQUA on more models in Table 8. Because robustness has little to do with the dataset, here we show the experimental results on the MMQA dataset as in the main text.
Table 8: This table shows the Rank and p-value of watermark image after some common transformations.
Models Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
LLaVA-NeXT Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
InternVL3 Rescale $1.03$ $4.85e^{-231}$ $1.36$ $4.85e^{-46}$
Rotate $1.07$ $9.17e^{-112}$ $1.61$ $1.59e^{-31}$
Gaussian $1.07$ $7.69e^{-229}$ $1.46$ $3.76e^{-39}$
Rescale + Rotate + Gaussian $1.06$ $6.04e^{-216}$ $1.79$ $8.24e^{-27}$
Qwen-VL-Chat Rescale $1.03$ $4.58e^{-154}$ $1.36$ $4.15e^{-45}$
Rotate $1.07$ $8.08e^{-106}$ $1.61$ $6.82e^{-32}$
Gaussian $1.07$ $5.26e^{-142}$ $1.46$ $9.14e^{-38}$
Rescale + Rotate + Gaussian $1.06$ $5.12e^{-125}$ $1.79$ $6.78e^{-25}$
Qwen2.5-VL-Instruct Rescale $1.03$ $3.19e^{-269}$ $1.36$ $7.68e^{-68}$
Rotate $1.07$ $7.54e^{-238}$ $1.61$ $4.15e^{-51}$
Gaussian $1.07$ $1.23e^{-251}$ $1.46$ $8.48e^{-62}$
Rescale + Rotate + Gaussian $1.06$ $5.47e^{-268}$ $1.79$ $6.49e^{-46}$

### v2

@@
-11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA
+arXiv:2506.10030v2[cs.CR] 28 Feb 2026
 Abstract
@@
 AQUA embeds semantic signals into synthetic images using two complementary methods: acronym-based triggers and spatial relationship cues.
-These techniques ensure watermark signals survive indirect watermark propagation from image retriever to textual generator, being efficient, effective and imperceptible.
+These techniques ensure watermark signals survive indirect watermark propagation from image retriever to textual generator, and are efficient, effective, and imperceptible.
 Experiments across diverse models and datasets show that AQUA enables robust, stealthy, and reliable copyright tracing, filling a key gap in Multimodal RAG protection.
+The implementation of AQUA is publicly available at https://github.com/tychenn/AQUA.
 1 Introduction
@@
 Data providers can issue watermark probe queries to RAG services.
-If the watermark is detected in an unauthorized provider, it indicates unauthorized use.
-Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks, but they often suffer from hallucinations and outdated knowledge learned in the static parameters.
-To mitigate these limitations, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020; Asai et al., 2023a) has emerged as a promising paradigm that augments LLMs with up-to-date external knowledge retrieved at inference time.
-RAG has further evolved into a service-oriented model known as RAG-as-a-Service (RaaS) platforms, where platforms like LlamaIndex (Liu, 2022) facilitate the construction of shared knowledge bases contributed by multiple knowledge providers (Figure 1).
-Importantly, these systems adopt a “usable but not visible” policy: RAG service provider can leverage the contributed knowledge without directly accessing the raw data.
-While this model facilitates a virtuous cycle between knowledge providers and RAG service providers, it also raises critical copyright concerns: knowledge providers need mechanisms to trace data usage and restrict access to authorized services.
-Since unauthorized RAG providers typically utilize the entire shared knowledge base, knowledge provider can embed watermarks at the knowledge base level.
-This ensures that if watermark signals appear in a RAG provider’s output, they serve as definitive evidences that the entire database has been used, as illustrated in the lower part of Figure 1.
+If the watermark is detected in an unauthorized RAG service, it indicates unauthorized use.
+Large Language Models (LLMs) often suffer from hallucinations and outdated knowledge, which Retrieval-Augmented Generation (RAG) mitigates by retrieving external knowledge at inference time (Lewis et al., 2020; Guu et al., 2020; Asai et al., 2023).
+RAG has further evolved into RAG-as-a-Service (RaaS), where platforms such as LlamaIndex (Liu, 2022) enable shared knowledge bases contributed by multiple providers (Figure 1).
+These systems follow a “usable but not visible” policy: service providers can use contributed knowledge without direct access to raw data.
+While RaaS enables a mutually beneficial ecosystem between knowledge providers and service platforms, it also introduces copyright and ownership challenges.
+In particular, providers require mechanisms to reliably trace data usage and restrict access to authorized services.
+Since unauthorized RAG providers typically utilize the entire shared knowledge base, a practical solution is to embed watermarks at the knowledge base level.
+The detection of these watermark signals in a provider’s output can then serve as strong evidence of unauthorized data usage (Figure 1, bottom).
+Existing watermarking methods for RaaS have primarily focused on textual knowledge (Jovanović et al., 2025; Guo et al., 2025).
+However, these methods are modality-specific, are limited to the text modality, and cannot be directly applied to non-textual knowledge due to the distinct characteristics of other modalities.
+In practice, RaaS systems increasingly integrate multimodal knowledge, combining textual and visual content (Riedler and Langer, 2024; Xia et al., 2024a; b).
+This creates a fundamental gap and leaves a critical vulnerability in the copyright protection of Multimodal RaaS.
+To address this gap, we focus on a representative subclass: text-to-text (T2T) Multimodal RAG, where the generator integrates retrieved image knowledge and the textual query to generate textual responses (Yasunaga et al., 2022; Chen et al., 2022; Lin and Byrne, 2022; Sun et al., 2024; Zhu et al., 2024).
 Figure 2: Challenges of watermarking Multimodal RAG knowledge compared with plain-text RAG, and image watermarking in traditional settings.
-Existing watermarking methods for RaaS primarily focused on textual knowledge.
-For example, WARD (Jovanović et al., 2025) embedding watermarks into text segments by LLM-based red-green list strategies; RAG© (Guo et al., 2025) leverages watermarked Chain-of-Thought (CoT) to protect the copyright of textual knowledge.
-However, these methods are modality-specific, limited to text modality and cannot be directly applied to non-textual knowledge due to the distinct characteristics of other modalities.
-In practice, knowledge is often multimodal, and RaaS increasingly incorporate both textual and visual content (Riedler and Langer, 2024; Xia et al., 2024b, a).
-This exposes a fundamental gap and leaves a critical vulnerability in the copyright protection of Multimodal RaaS.
-To address this gap, we focus on a representative subclass: text-to-text (T2T) Multimodal RAG, where generator integrates retrieved image knowledge and textual query to generate textual responses (Yasunaga et al., 2022; Chen et al., 2022a; Lin and Byrne, 2022; Sun et al., 2024; Zhu et al., 2024).
 Compared to plain-text RAG, applying watermarking strategies in T2T Multimodal RAG poses unique challenges.
-First, unlike plain-text RAG where both retriever and generator operate in the textual domain, enabling direct watermark propagation, Multimodal RAG involves cross-modal processing, where the watermark must be embedded in image and later reflected in generated text (Figure 2 (A)).
-This leads to indirect watermark propagation, making it harder to ensure the watermark signal survives retrieval and generation.
-Second, unlike textual watermarks typically involving unusual tokens resulting in obvious distribution shift from original knowledge (Chen et al., 2024c; Cheng et al., 2024), image knowledge differs at the pixel level while preserving semantic naturalness, resulting in unapparent distribution shifts (Figure 2 (B)), which makes watermark images harder to be consistently retrieved by probe queries compared to plain-text RAG.
-Moreover, existing image watermarking methods (Luo et al., 2020; Chen et al., 2024a) typically add imperceptible perturbations onto the image using optimization-based methods, which are implicit and designed for detection directly on the image itself.
-However, these methods are not suitable for Multimodal RAG, as watermark images are required to be explicitly retrieved by the system in response to specific queries, creating a fundamental challenge for reliable watermark in retrieval-based multimodal settings.
-To address above challenges in image knowledge copyright protection, we propose AQUA, a novel watermarking framework tailored for T2T Multimodal RAG.
-Specifically, AQUA watermarking framework includes two complementary watermarking methods: $\emph{AQUA}_{acronym}$ and $\emph{AQUA}_{spatial}$.
-$\textit{AQUA}_{acronym}$ addresses indirect watermark propagation by embedding uncommon acronyms and their full names into synthetic images.
-In the verification phase, these acronyms are decoded through the Optical Character Recognition (OCR) abilities of generators (Vision-Language Models (VLMs) (Achiam et al., 2023; Team et al., 2023; Huang et al., 2023) ) to generate detectable textual response: the full name of the acronyms.
+First, while text-based RAG supports direct watermark propagation, Multimodal RAG requires embedding the watermark in images and reflecting it in generated text, resulting in indirect propagation that is harder to preserve.
+Second, unlike textual watermarks typically involving unusual tokens resulting in obvious distribution shift from original knowledge (Chen et al., 2024c; Cheng et al., 2024), image knowledge differs only at the pixel level while preserving semantic naturalness, resulting in unapparent distribution shifts (Figure 2 (B)), that reduce retrievability.
+Moreover, existing image watermarking methods (Luo et al., 2020; Chen et al., 2024a) rely on implicit perturbations designed for image-level detection, but Multimodal RAG requires watermarks to be explicitly retrieved through queries, making them unsuitable for retrieval-based multimodal settings.
+To address the above challenges in image knowledge copyright protection, we propose AQUA, a novel watermarking framework tailored for T2T Multimodal RAG.
+Specifically, the AQUA watermarking framework includes two complementary watermarking methods: $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$.
+$\textbf{AQUA}_{acronym}$ addresses indirect watermark propagation by embedding uncommon acronyms and their full names into synthetic images.
+In the verification phase, these acronyms are decoded through the Optical Character Recognition (OCR) abilities of Vision-Language Models (VLMs) (Achiam et al., 2023; Gemini Team et al., 2023; Huang et al., 2023) to generate detectable textual responses: the full name of the acronyms.
 Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing.
-For models with limited OCR ability, $\emph{AQUA}_{spatial}$ is designed to create synthetic images with special object configurations (e.g.
-unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries.
+For models with limited OCR ability, $\textbf{AQUA}_{spatial}$ is designed to create synthetic images with special object configurations (e.g., uncommon positional relationships), and to leverage generators’ understanding of spatial semantics to answer position-related probe queries.
 These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator.
-Both methods introduces semantic distinctiveness by embedding subtle semantic cues into natural-looking images, allowing explicit retrieval while maintaining a high retrieval rate.
+Both methods introduce semantic distinctiveness by embedding subtle semantic cues into natural-looking images, allowing explicit retrieval while maintaining a high retrieval rate.
 Together, these two methods provide a flexible, robust solution to the unique challenges of watermarking in Multimodal RAG systems, supporting both black-box and white-box deployments.
-Despite simplicity, our novel insights of using synthetic images with special acronyms texts and special positional relationships as watermark carriers are particularly effective and efficient in bridging the gap between image-based watermarking and textually detectable outputs, enabling robust copyright tracing in Multimodal RAG.
-We evaluate AQUA across diverse Multimodal RAG and datasets spanning different domains.
-The experimental results demonstrate that AQUA (1) enables the watermark images to be retrieved and reflected in the generated textual output, (2) prevent false positives from common image content, (3) remain imperceptible to users and undetectable by unauthorized filtering mechanisms, and (4) is robust to common image transformations.
-Our contribution can be summarized as follow:
-  • We propose AQUA, the first watermarking framework tailored for image knowledge copyright protection in Multimodal RAG systems, addressing indirect watermark propagation, and successful retrieval under unapparent distribution shifts and explicit watermark injection;
-  • We design two complementary watermarking strategies, $\emph{AQUA}_{acronym}$, $\emph{AQUA}_{spatial}$ to support more realistic black-box scenarios;
-  • We perform comprehensive experiments on two well-known datasets, utilizing four prevalent pretrained VLMs (LLaVA-NeXT, InternVL3, Qwen-VL-Chat and Qwen2.5-VL-Instruct) to assess the effectiveness, harmlessness, stealthiness and robustness of AQUA.
+Despite its simplicity, our novel insights of using synthetic images with special acronyms text and special positional relationships as watermark carriers are particularly effective and efficient in bridging the gap between image-based watermarking and textually detectable outputs, enabling robust copyright tracing in Multimodal RAG.
+We evaluate AQUA across diverse Multimodal RAG systems and datasets spanning different domains.
+The experimental results demonstrate that AQUA (1) enables the watermark images to be retrieved and reflected in the generated textual output, (2) prevents false positive retrieval from common image content, (3) remains imperceptible to users and undetectable by unauthorized filtering mechanisms, and (4) is robust to attacks such as image transformations and regeneration.
+Our contribution can be summarized as follows:
+  • We propose AQUA, the first watermarking framework for image knowledge copyright protection in Multimodal RAG, addressing indirect watermark propagation and successful retrieval under unapparent distribution shifts and explicit watermark injection.
+  • We design two complementary watermarking strategies, $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$, to support more realistic black-box scenarios.
+  • We conduct comprehensive experiments on two RAG datasets and RAG architectures to demonstrate the effectiveness, harmlessness, stealthiness and robustness of AQUA.
   • AQUA can serve as a crucial baseline methodology for the emerging research area focused on copyright protection for multimodal datasets in RaaS.
-2 Related Works
+2 Related Work
 2.1 Multimodal Retrieval-Augmented Generation
-Plain-text Retrieval-Augmented Generation (RAG).
-(Lewis et al., 2020; Singh et al., 2021; Wang et al., 2023; Asai et al., 2023b; Xu et al., 2024; Zhao et al., 2024; Tan et al., 2025) proposed several RAG frameworks to address the hallucination in LLMs.
-They mainly focus on textual external knowledge, and these methods are inadequate to process and integrate non-textual modalities inherent in real-world information.
-Multimodal Retrieval-Augmented Generation (Multimodal RAG).
-(Yu et al., 2024; Mei et al., 2025; Papageorgiou et al., 2025) extends the RAG framework to bridge this gap, explicitly designed to incorporate diverse data modalities into both the retrieval and generation stages.
-A common strategy for enabling cross-modal retrieval is to employ powerful multimodal encoders (e.g.
-CLIP (Radford et al., 2021) ), to map different modalities (e.g., text and images) into a shared semantic embedding space.
+Relying only on textual information is a limited approach for describing the intricate nature of the physical world.
+Yu et al.
+(2024); Mei et al.
+(2025); Papageorgiou et al.
+(2025) extend the text-only RAG framework to multimodal settings and explicitly incorporate diverse data modalities into both the retrieval and generation stages.
+A common strategy for enabling cross-modal retrieval is to employ powerful multimodal encoders (e.g., CLIP (Radford et al., 2021)), to map different modalities (e.g., text and images) into a shared semantic embedding space.
 This unification allows standard vector search algorithms like cosine similarity to retrieve relevant items across modalities based on semantic relatedness.
@@
 Several watermarking approaches have been proposed to protect the copyright of textual knowledge in RAG.
-WARD (Jovanović et al., 2025) uses the LLM red-green list watermarking technology to watermark all the texts in the RAG knowledge base (Kirchenbauer et al., 2023; Gloaguen et al., 2024) .
+WARD (Jovanović et al., 2025) uses the LLM red-green list watermarking technology to watermark all text in the RAG knowledge base (Kirchenbauer et al., 2023; Gloaguen et al., 2024).
 RAG-WM (Lv et al., 2025) presents a black-box RAG watermarking approach that leverages interactions among multiple LLMs to generate high-quality watermarks.
@@
 DMI-RAG (Liu et al., 2025) performs dataset membership inference by injecting a small number of synthetic, watermarked "canary" documents into the Intellectual Property (IP) dataset.
-However, existing methods on watermarking knowledge base in RAG system have exclusively focused on purely textual data.
+However, existing methods on watermarking knowledge bases in RAG systems have exclusively focused on purely textual data.
 To the best of our knowledge, no prior work has addressed the protection of knowledge copyright in Multimodal RAG systems, particularly those integrating image and text modalities, via watermarking techniques.
 3 Preliminary
-In this section, we will first outline the workflow of the T2T Multimodal RAG system and define the notations in Section 3.1.
-Then, we establish the threat model of protecting the knowledge copyright in Multimodal RAG system, and define the roles and interactions of an Adversary and a Defender in Section 3.2.
+In this section, we outline the workflow of the T2T Multimodal RAG system and define the notation in Section 3.1.
+Then, we establish the threat model for knowledge copyright protection in Multimodal RAG systems in Section 3.2.
 3.1 Multimodal RAG System Workflow
-The T2T Multimodal RAG system contains three components: a retriever $\mathcal{E}$, a generator $\mathcal{G}$, and an external image knowledge base $D$.
-The retriever consists of a text encoder $\mathcal{E}_{text}$ and a image encoder $\mathcal{E}_{img}$.
-Images $I_{i}$ in the external knowledge base $D=\{I_{1},\dots,I_{n}\}$ are pre-processed to a latent space through the image encoder: $e_{I_{i}}=\mathcal{E}_{img}(I_{i})\in\mathbb{R}^{d}$.
-Knowledge Retrieval.
-The retriever accepts the user’s text query $T$ as input, and process it into the same latent space as image: $e_{T}$: $e_{T}=\mathcal{E}_{text}(T)\in\mathbb{R}^{d}$.
-Then the retriever employs a similarity function, $\text{Sim}(\cdot,\cdot):=\mathbb{R}^{d}\times\mathbb{R}^{d}\to\text{Score}$ (e.g., cosine similarity), to find the most relevant image knowledge according to user’s text query: $s_{i}=\text{Sim}(e_{T},e_{I_{i}})$.
+The T2T Multimodal RAG system consists of three components: a retriever $\mathcal{E}$, a generator $\mathcal{G}$, and an external image knowledge base $D$.
+The retriever includes a text encoder $\mathcal{E}_{text}$ and an image encoder $\mathcal{E}_{img}$.
+Images $I_{i}$ in the external knowledge base $D=\{I_{1},\dots,I_{n}\}$ are preprocessed to a latent space through the image encoder: $e_{I_{i}}=\mathcal{E}_{img}(I_{i})\in\mathbb{R}^{d}$.
+The retriever accepts the user’s text query $T$ as input, and processes it into the same latent space as images $e_{T}$$\mathcal{E}_{text}(T)\in\mathbb{R}^{d}$.
+Then the retriever employs a similarity function, $\text{Sim}(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\text{Score}$ (e.g., cosine similarity), to find the most relevant image knowledge according to the user’s text query: $s_{i}=\text{Sim}(e_{T},e_{I_{i}})$.
 Based on these similarity scores $s_{i}$, the retriever selects the top-k most relevant images as output:
 $D_{retrieved}=\mathcal{R}(D,T,k)=\{I_{s_{(1)}},I_{s_{(2)}},\dots,I_{s_{(k)}}\},\;\text{where }S_{\text{top-k}}=\{s_{(1)},s_{(2)},\dots,s_{(k)}\}$ (1)
-Augmented Generation.
 The original text query $T$ and the retrieved set of images $D_{retrieved}$ are combined and passed to the generator $\mathcal{G}$ to produce the final answer: $A=\mathcal{G}(D_{retrieved},T)$
 3.2 Threat Model
-We consider the image knowledge copyright protection in Multimodal RAG service.
-Defender represents the knowledge provider, aiming to detect and prevent unauthorized use of their proprietary image knowledge by external Multimodal RAG services.
+We consider image knowledge copyright protection in Multimodal RAG services.
+Defender represents the knowledge provider, aiming to detect and audit unauthorized use of its proprietary image knowledge by external Multimodal RAG services.
 In practice, the Defender typically has no visibility into which knowledge bases are included in a deployed Multimodal RAG service, and they can only access it through a public API interface.
-Defender can only operate on their own datasets to implement protection mechanisms such as injecting watermarks before contributing their data to a RaaS.
-Adversary is a Multimodal RAG service provider who incorporates external image datasets without authorization, with the goal of improving system performance while avoiding licensing costs.
-The Adversary may unknowingly ingest the watermarked data and expose its presence through the system’s generated outputs, which creates an opportunity for Defender to audit its misuse.
+Defender can only operate on its own datasets to implement protection mechanisms such as injecting watermarks before contributing its data to a RaaS.
+Adversary is a Multimodal RAG service provider who incorporates image datasets without authorization, with the goal of improving system performance while avoiding licensing costs.
+Adversary may unknowingly ingest the watermarked data and expose its presence through the system’s generated outputs, creating an opportunity for Defender to audit its misuse.
 4 Methodology
-AQUA is a watermarking framework designed to protect the image knowledge copyrights in Multimodal RAG service, meeting four key requirements: effectiveness, harmlessness, stealthiness, and robustness.
-In this section, we instantiate the AQUA framework with two complementary watermarking methods, $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$.
-For each method, we will first introduce the principle of designing watermarks and then clarify how this method can be verified in statistical strategies.
-Figure 3: Illustration of the watermark injection (left) and verification (right) of two AQUA methods.
-4.1 $\textit{AQUA}_{arconym}$
+AQUA is a watermarking framework designed for copyright protection of image knowledge in Multimodal RAG services, meeting four key requirements: effectiveness, harmlessness, stealthiness, and robustness.
+In this section, we instantiate the AQUA framework with two complementary watermarking methods, $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$.
+Figure 3: Illustration of the watermark injection (left) and verification (right) of AQUA.
+4.1 $\textbf{AQUA}_{acronym}$
 Watermark Injection.
-$\textit{AQUA}_{acronym}$ addresses indirect watermark propagation from image knowledge to detectable textual output by embedding uncommon acronyms and their full names into synthetic images.
+$\textbf{AQUA}_{acronym}$ addresses indirect watermark propagation from image knowledge to detectable textual output by embedding uncommon acronyms and their full names into synthetic images.
 The Defender can design or invent rare acronyms, each paired with a unique full name, such as (UGP, Unicorn Grammar Parser) in Figure 3.
@@
 Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing.
-The acronym pair can also be generated in large quantities using LLM (e.g., Gemini-2.5-Pro), with the ability of In-Context Learning (ICL) (Brown et al., 2020) and the prompt provided in Appendix A.1, and more examples are relegated in Appendix A.2.
+Acronym pairs can also be generated in large quantities using an LLM (e.g., Gemini-2.5-Pro), via In-Context Learning (ICL) (Brown et al., 2020) and the prompt provided in Appendix A.1, and more examples are provided in Appendix A.2.
 Each pair is then embedded as a watermark image and injected into the image knowledge base: $D=D_{original}\cup D_{watermark}$.
@@
 Watermark Verification.
-In verification phase, these acronyms are decoded through the OCR ability of generator, to generate detectable textual responses: the full name of the acronyms.
+In the verification phase, these acronyms are decoded through the generator’s OCR ability, producing detectable textual responses: the full name of the acronyms.
 Each watermark image has its own probe query $T_{probe}$ which can be used by the knowledge provider to detect unauthorized use.
 The $T_{probe}$ consists of two parts: a trigger $T_{trigger}$, used by the retriever to retrieve the watermark images, and an instruction $T_{instruction}$, which prompts the generator to generate the watermark-included responses that can be detected.
-We can formulate this construction as: $T_{probe}=T_{tigger}\oplus T_{instruction}$.
-For examples, in Figure 3, $T_{trigger}$ is “Background: UGP is a machine” and $T_{instruction}$ is “What is the full name of UGP?”.
-Following (Wu et al., 2024; Ha et al., 2025) to verify the watermark signal, we define a strict exact match protocol $\text{Eval}(\cdot,\cdot)$ based on a normalization function $\text{Norm}(\cdot)$ that lowercases and strips whitespace from both generated output $O_{RAG}$ and the verification signature $S$:
+We can formulate this construction as $T_{probe}=T_{trigger}\oplus T_{instruction}$.
+For example, in Figure 3, $T_{trigger}$ is “Background: UGP is a machine” and $T_{instruction}$ is “What is the full name of UGP?”.
+To verify the watermark signal, we define a strict substring-match protocol $\text{Eval}(\cdot,\cdot)$ based on a normalization function $\text{Norm}(\cdot)$ that lowercases and strips whitespace from both generated output $O_{RAG}$ and the verification signature $S$:
 $\text{Eval}(O_{RAG},S)=\mathbb{I}[\text{Norm}(S)\subseteq\text{Norm}(O_{RAG})]$ (2)
-where $\mathbb{I}[\cdot]$ is the indicator function, returning $1$ if the condition (substring presence) is true, and 00 otherwise.
+where $\mathbb{I}[\cdot]$ is the indicator function, returning $1$ if the condition (substring presence) is true, and 0 otherwise.
 The predefined signature (e.g., "Unicorn Grammar Parser") serves as the ground truth.
-Due to the inherent randomness of generation (e.g., temperature, top-k/top-p sampling) (Ackley et al., 1985; Fan et al., 2018; Holtzman et al., 2019), the presence of a watermark signal is not guaranteed even when the corresponding image is retrieved.
+Due to the inherent randomness of generation (e.g., temperature, top-k/top-p sampling) (Fan et al., 2018; Holtzman et al., 2019), the presence of a watermark signal is not guaranteed even when the corresponding image is retrieved.
 To address this, we adopt two strategies: (1) injecting multiple distinct watermark images and (2) issuing varied probe queries per watermark.
 We define the Verification Success Rate (VSR) as:
-$\text{VSR}=\frac{1}{N_{wm}\cdot N_{ds}}\sum_{j=1}^{N_{wm}}\sum_{i=1}^{N_{ds}}\text{Eval}_{j}(O_{{RAG}_{i}},S_{i})$ (3)
-where $N_{wm}$ is the number of watermark images and $N_{ds}$ is the number of distinct queries per image.
-$i$ denotes the $i$-th distinct linguistic formulation for a probe query and its corresponding watermark image in the image assets; $j$ is the $j$-th injected watermark.
+$\text{VSR}=\frac{1}{N_{wm}\cdot N_{ds}}\sum_{j=1}^{N_{wm}}\sum_{i=1}^{N_{ds}}\text{Eval}(O_{{RAG}_{i,j}},S_{i})$ (3)
+where $N_{wm}$ is the number of watermark images and $N_{ds}$ is the number of distinct queries per watermark image.
+$j$ denotes the $j$-th injected watermark, and $i$ denotes the $i$-th probe query with its corresponding signature $S_{i}$.
+$O_{{RAG}_{i,j}}$ is the generated output when the $i$-th probe query is issued against the $j$-th watermark image.
 Hypothesis Testing.
 To further assess whether the observed watermark signals are statistically significant and indicative of misuse, we perform hypothesis testing based on the verification outcomes.
-Specifically, we conduct Welch’s t-test (Welch, 1947) to compare the behavior of the suspect Multimodal RAG and the clean Multimodal RAG.
-Null Hypothesis ($\mathcal{H}_{0}$) indicates there is no statistical evidence suggesting the suspect Multimodal RAG including the watermark image datasets: $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, where the VSR of the suspect Multimodal RAG is equal to the VSR of the clean one.
-Using the sample means, variances, counts, and approximated degrees of freedom via the Welch-Satterthwaite equation (Satterthwaite, 1941, 1946), we compute the t-statistic.
+Specifically, following Xu et al.
+(2023), we conduct Welch’s t-test (Welch, 1947) to compare the behavior of the suspect Multimodal RAG and the clean Multimodal RAG.
+The null hypothesis ($\mathcal{H}_{0}$) indicates that there is no statistical evidence suggesting the suspect Multimodal RAG includes the watermark image datasets: $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, where the VSR of the suspect Multimodal RAG is equal to the VSR of the clean one.
+Using the sample means, variances, counts, and approximated degrees of freedom via the Welch-Satterthwaite equation (Satterthwaite, 1941; 1946), we compute the t-statistic.
 The p-value is compared against a significance level (e.g., $\alpha=0.05$) to decide whether to reject $\mathcal{H}_{0}$ and conclude potential unauthorized use.
-4.2 $\textit{AQUA}_{spatial}$
+The practical deployment of AQUA presents unique challenges due to the specific state of each target RAG database; these issues are discussed in detail in Appendix B.1.
+4.2 $\textbf{AQUA}_{spatial}$
 Watermark Injection.
-For those models with limited OCR capabilities, we propose $\textit{AQUA}_{spatial}$, which is designed to create synthetic images with special object configurations (e.g., unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries.
+For models with limited OCR capabilities, we propose $\textbf{AQUA}_{spatial}$, which is designed to create synthetic images with special object configurations (e.g., unusual positional relationships), and to leverage generators’ understanding of spatial semantics to answer position-related probe queries.
 Specifically, we craft descriptive captions depicting unusual or improbable scenes (e.g., “A red apple on the head of a reading dog.”) and generate corresponding images using a diffusion model (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022).
+To ensure semantic uniqueness, these captions are constructed from low co-occurrence concept pairs and further filtered by language model perplexity to balance distinctiveness with naturalness (Appendix A.4).
 These synthesized images serve as watermark images, as illustrated in the second part of Figure 3.
-Similar to $\textit{AQUA}_{acronym}$, these watermark images are injected into the dataset and can be scaled using LLM-based in-context generation of diverse captions.
-More examples are relegated to Appendix A.2.
-These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator.
+Similar to $\textbf{AQUA}_{acronym}$, these watermark images are injected into the dataset and can be scaled using LLM-based in-context generation of diverse captions.
+More examples and image caption templates are provided in Appendix A.2 and Appendix A.4, respectively.
 Watermark Verification and Hypothesis Testing.
-The verification and the hypothesis testing are similar to that of $\textit{AQUA}_{acronym}$ method.
+The verification and the hypothesis testing are similar to those of the $\textbf{AQUA}_{acronym}$ method.
 Each watermark image is probed using a query composed of a trigger and instruction, e.g., $T_{trigger}$ = “There is a dog reading a book.” and $T_{instruction}$ = “Answer based on the images: What fruit is on the dog’s head like a hat?”.
 The expected signature is “Apple”.
-As before, the system output is evaluated using the exact-match protocol $\text{Eval}(\cdot,\cdot)$, and Welch’s t-test is applied to determine whether the suspect system statistically includes the watermarked dataset.
+As before, the system output is evaluated using the substring-match protocol $\text{Eval}(\cdot,\cdot)$, and Welch’s t-test is applied to determine whether the suspect system statistically includes the watermarked dataset.
 4.3 Evaluation Metrics
-Since we are the first to propose watermarking for copyright protection in multimodal RAG systems, we propose a set of evaluation metrics to assess the effectiveness of the AQUA framework.
-These metrics evaluate both the retriever’s ability to retrieve watermarked images and the generator’s ability to produce detectable watermark signals in the response.
+We evaluate AQUA using multiple metrics that capture both retrieval and generation performance.
+Verification success rate and hypothesis-testing-based assessments quantify the overall effectiveness of watermark detection.
+In addition, we introduce Rank and Conditional Generation Success Rate (CGSR).
 Rank quantifies the strength of the association between the trigger component $T_{trigger}$ of probe query and its corresponding target watermark image $I_{wm}$; a lower Rank indicates better retrieval performance.
-For a given query, $D_{retrieved}=(I_{1},I_{2},\dots,I_{k})$ indicates the top-$k$ retrieved images knowledge.
+For a given query, $D_{retrieved}=(I_{1},I_{2},\dots,I_{k})$ indicates the top-$k$ retrieved images.
 The Rank is defined as the 1-based index $r$ of $I_{wm}$ within $D_{retrieved}$.
-If $I_{wm}$ is not present within the top $k$ retrieved images, a penalty value, set to twice the retrieval depth ($2k$), is assigned.
+If $I_{wm}$ is not present within the top-$k$ retrieved images, a penalty value, set to twice the retrieval depth ($2k$), is assigned.
 Formally, the Rank is calculated as:
@@
 Conditional Generation Success Rate (CGSR) measures the proportion of successful generations where the verification signature $S$ is correctly produced, given that the corresponding watermark image has been successfully retrieved.
-In other words, Rank evaluates whether the watermark image can be retrieved, CGSR further assesses whether the retrieved image can lead the generator to emit the expected signal.
 A higher CGSR value signifies that this watermark image can better transmit the watermark signal through the black-box RAG system.
-Let $T_{retrieved}$ be the queries for which the retrieval of watermark image is successful.
+Let $T_{retrieved}$ be the queries for which the retrieval of the watermark image is successful.
 The CGSR is then defined as the success rate over the subset of successful retrievals:
 $\text{CGSR}=\frac{\sum_{t\in T_{retrieved}}\text{Eval}(O_{RAG}^{(t)},S^{(t)})}{|T_{retrieved}|}$ (5)
-SimScore quantifies the output quantifies the semantic similarity between a watermark probe query and a benign query with similar intent, as judged by an LLM (Gemini-2.5-Pro), with scores ranging from 0 to 100%.
-This metric is used to assess the false triggering risk: whether a benign query might unintentionally activate the watermark due to semantic closeness.
-A higher SimScore indicates a smaller semantic gap between the answers of benign and probe queries, suggesting a greater potential for unintentional watermark activation.
-The prompting details are provided in Appendix A.1.
+SimScore quantifies the semantic similarity between two textual strings, as judged by an LLM (Gemini-2.5-Pro), with scores ranging from 0 to 100%.
+In our evaluation, we apply SimScore to query pairs to assess false triggering risk and to response pairs to measure whether the watermark distorts normal system outputs.
 5 Experiments
@@
 Both datasets contain a large number of QA pairs, and the questions can only be answered by combining knowledge of modalities such as text, images, and tables.
-Our intention is to use AQUA to protect the copy right of image modality.
 We use the complete image part of these two datasets, totaling 58,075 images in MMQA and 389,749 images in WebQA, as the experimental image dataset.
-[…]
-Devices. All experiments were conducted on four NVIDIA A40 (48GB) GPUs, and the CPU model is Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz.
-5.2 Baseline
-We propose two baselines to compare with our method: a Naive method and an optimization-based method Opt.
-Naive baseline uses usual images as watermark images.
-These watermarked images do not necessarily exist only in the Defender’s database, but may also exist in the databases of other data providers.
-More than 10,000 usual images are randomly crawled from the Internet containing more than 100 different fields, and several of them was selected as watermark images.
-Opt.
-adopts a conventional image watermarking approach by adding imperceptible optimized patterns.
-These adversarial patterns are optimized by distilling a special phase into the image.
-Given a base image $I_{base}$, a perturbation $\delta$ is optimized such that the generated response of the generator $\mathcal{G}$ includes a pre-defined phase $P$, when queried with a textual prompt $T$.
-The objective can be formulated as minimizing the cross-entropy loss between the generated response and the target signature $S$ :
-$\min_{\delta}L(\mathcal{G}(I_{base}+\delta,T),P)\quad$ (6)
+Multimodal RAG Components.
+We use the Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021), specifically the openai/clip-vit-large-patch14 variant as Retriever.
+Cosine Similarity is used to compute the similarity between text and image.
+Following the usual search strategies (Caffagni et al., 2024; Mortaheb et al., 2025; Ha et al., 2025), we set clip-top-k=5, ensuring the retriever selects the five most relevant images as context.
+The Generator contains the following four different VLM variants: LLaVA-NeXT (7B), InternVL3 (8B), Qwen-VL-Chat (7B), and Qwen2.5-VL-Instruct (7B) (Liu et al., 2024; Chen et al., 2024d; Bai et al., 2023; Bai and others, 2025a).
+To control the diversity of the outputs, we configure the decoding process for each VLM using standard sampling parameters: sampling temperature ($\text{T}=1.2$), top-k sampling ($\text{top\_k}=5$), nucleus sampling ($\text{top\_p}=0.9$).
+These settings are maintained consistently across experiments.
+5.2 Baselines
+We propose two baselines to compare with AQUA: a Naive random select method and an optimization-based method.
+Naive baseline uses common images as watermark images.
+These images are not unique to the Defender’s database but may also appear in databases of other data providers.
+Specifically, we randomly crawled more than 10,000 images from the Internet over 100 domains, and selected a subset of them as watermark images.
+The optimization-based method follows a conventional image watermarking approach by embedding imperceptible optimized patterns.
+These adversarial patterns are optimized by distilling a special phrase into the image.
+Specifically, a perturbation $\delta$ is optimized and added to a base image $I_{base}$ such that, when the watermarked image is queried with a textual prompt $T_{probe}$, the generator $\mathcal{G}$ produces an output containing a predefined signature $S$.
+The optimization objective is to minimize the cross-entropy loss between the generated response and the target signature:
+$\min_{\delta}L(\mathcal{G}(I_{base}+\delta,T_{probe}),S)\quad$ (6)
 We adopt Projected Gradient Descent (PGD) (Goldstein, 1964; Levitin and Polyak, 1966) to optimize the perturbation iteratively, as it is a widely-adopted and effective adversarial perturbation generation method:
-$\delta_{t+1}=\Pi_{\|\cdot\|_{p}\leq\epsilon}\left(\delta_{t}-\alpha\cdot\nabla_{\delta_{t}}L(\mathcal{M}(I_{base}+\delta_{t},q),P)\right)$ (7)
-where $\alpha$ represents the step size (learning rate), and projection operator $\Pi_{\|\cdot\|_{p}\leq\epsilon}(\cdot)$ ensures the perturbation remains within an $L_{p}$-norm ball of radius $\epsilon$, preserving visual imperceptibility.
+$\delta_{t+1}=\Pi_{\|\cdot\|_{p}\leq\epsilon}\left(\delta_{t}-\alpha\cdot\nabla_{\delta_{t}}L(\mathcal{G}(I_{base}+\delta_{t},T_{probe}),S)\right)$ (7)
+where $\alpha$ represents the step size (learning rate), and the projection operator $\Pi_{\|\cdot\|_{p}\leq\epsilon}(\cdot)$ ensures the perturbation remains within an $L_{p}$-norm ball of radius $\epsilon$, preserving visual imperceptibility.
 The final watermarked image is $I_{wm}=I_{base}+\delta^{*}$.
-5.3 Effectiveness of AQUA
 Table 1: Effectiveness of AQUA.
 Models indicate which model is used as the generator.
-$\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ represent the two watermarking methods.
+$\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ represent the two watermarking methods.
 Naive and Opt.
-denotes the baseline methods.
+denote the baseline methods.
 Models Methods MMQA WebQA
@@
 LLaVA - NeXT Naive $2.86$ $28.16\%$ $0.32$ $4.56$ $13.28\%$ $0.93$
-Opt. $1.45$ $31.03\%$ $3.33e^{-4}$ $1.90$ $22.86\%$ $3.94e^{-2}$
+Opt. $1.45$ $1.03\% 3.3$$33e^{-4}$ $1.90$ $22.86\%$ $3.94e^{-2}$
+$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $\bm{85.36\%}$ $\bm{0.0}$ $\bm{1.05}$ $78.73\%$ $\bm{9.47e^{-182}}$
+$\textbf{AQUA}_{spatial}$ $1.29$ $75.38\%$ $1.07e^{-67}$ $1.85$ $\bm{86.45\%}$ $2.3e^{-45}$
+InternVL3 Naive $2.86$ $27.11\%$ $0.41$ $4.56$ $17.12\%$ $0.65$
+Opt. $1.45$ $19.34\%$ $5.39e^{-3}$ $1.90$ $19.45\%$ $3.87e^{-3}$
+$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $\bm{85.11\%}$ $\bm{6.29e^{-289}}$ $\bm{1.05}$ $\bm{78.34\%}$ $\bm{2.88e^{-129}}$
+$\textbf{AQUA}_{spatial}$ $1.29$ $75.72\%$ $1.49e^{-50}$ $1.85$ $72.46\%$ $4.31e^{-26}$
+Qwen-VL -Chat Naive $2.86$ $15.79\%$ $0.59$ $4.56$ $5.71\%$ $0.91$
+Opt. $1.45$ $21.29\%$ $9.05e^{-3}$ $1.90$ $18.91\%$ $1.21e^{-3}$
+$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $75.28\%$ $\bm{1.05e^{-162}}$ $\bm{1.05}$ $\bm{77.86\%}$ $\bm{1.24e^{-128}}$
+$\textbf{AQUA}_{spatial}$ $1.29$ $\bm{78.92\%}$ $1.35e^{-60}$ $1.85$ $68.46\%$ $9.63e^{-35}$
+Qwen2.5- VL-Instruct Naive $2.86$ $38.15\%$ $0.25$ $4.56$ $15.87\%$ $0.86$
+Opt. $1.45$ $19.96\%$ $7.35e^{-3}$ $1.90$ $18.51\%$ $6.77e^{-3}$
+$\textbf{AQUA}_{acronym}$ $\bm{1.03}$ $\bm{99.61\%}$ $\bm{0.0}$ $\bm{1.05}$ $\bm{96.68\%}$ $\bm{6.6e^{-145}}$
+$\textbf{AQUA}_{spatial}$ $1.29$ $98.42\%$ $8.29e^{-72}$ $1.85$ $89.85\%$ $2.92e^{-49}$
+5.3 Effectiveness of AQUA
+This section presents an empirical evaluation of the effectiveness of the proposed AQUA framework.
+Performance is quantified using Rank and CGSR metrics, with results summarized in Table 1.
+Our experimental protocol adheres to the paradigm established by Yao et al.
+(2024), utilizing a dataset of 50 distinct watermark images.
+Each image is subjected to 10 unique probe queries with diverse syntactic structures.
+To ensure statistical robustness, the entire experiment is replicated 10 times.
+Welch’s t-test is conducted to assess the statistical significance of the detection results.
+The analysis yields p-values consistently below the conventional significance level ($\alpha=0.05$), which leads to the rejection of the null hypothesis, $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$.
+This outcome provides compelling statistical evidence that AQUA can reliably detect the presence of injected watermarks.
+For a complementary analysis, and in accordance with the methodology of Jovanović et al.
+(2025), Appendix C.1 provides the results of a Two-proportion Z-test, along with experimental results across models of varying parameter sizes.
+Analysis of Query Efficiency.
+Although the optimization-based method can also ultimately achieve a statistically significant result (i.e., a low p-value) to reject the null hypothesis, the number of queries required to do so is a critical performance metric in real-world applications, particularly where queries are costly or limited.
+We therefore evaluate query efficiency by measuring the number of queries each method needs to reach the significance threshold.
+As depicted in Figure 4(a), both $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ achieve a p-value below the significance level within 30 queries.
+In contrast, the Opt.
+baseline requires over 200 queries to attain the same level of statistical confidence.
+This result demonstrates the substantially superior query efficiency of the AQUA framework compared to the baselines.
+(a) p-value vs. query times
+(b) TPR vs. FPR
+Figure 4: Diagnostics of AQUA watermark detection.
+FPR vs.
+TPR.
+To further validate the effectiveness of AQUA, we analyze its True Positive Rate (TPR) against its False Positive Rate (FPR), as shown in Figure 4(b).
+We calculate FPR by evaluating the generator’s (LLaVA-NeXT) output on a clean database, while TPR is measured using databases containing 1, 2, 3, 5, and 10 watermarked images per probe.
+The substantial distance of the AQUA curve from the random baseline indicates a strong statistical separation between watermarked and clean distributions.
+This characteristic is crucial, as it confirms that AQUA can achieve a high detection rate while keeping the false positive rate exceptionally low, thus validating the method’s precision and reliability.
+5.4 Harmlessness of AQUA
+Normal Query.
+To verify the harmlessness of our watermark, we evaluated the system’s responses to over 10,000 benign queries sourced from the MMQA and WebQA datasets.
+A watermark is considered harmless if it is neither retrieved nor reflected in the generated output during the system’s normal operation.
+In our experiments, with a single watermarked image embedded in the knowledge base, the retrieval rate for the watermarked content is $0\%$ for both the $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ variants.
+Concurrently, the CGSR is also $0\%$ across all four generators tested.
+These results confirm that our verification signature remains latent during standard interactions and does not interfere with the generation of correct responses to benign queries.
+Table 2: Examples of relevant queries and corresponding results.
+Type Example Probe Query Example Relevant Query Rank SimScore $\uparrow$
+Acronym- replace What is the subtitle of UGP? What is the subtitle of ATM? $10.00$ $100\%$
+Acronym- no_instru What is the subtitle of UGP? What is UGP? $1.07$ $70.18\%$
+Spatial- imprecise What fruit is the monkey holding like a phone? What is the monkey holding? $2.93$ $75.87\%$
+Relevant Queries.
+We employ relevant queries to assess whether the embedded watermark interferes with the model’s standard output when processing prompts highly similar to the probe query.
+The experimental results on LLaVA-NeXT and MMQA (Table 2) indicate that if the unique acronym within the probe query is substituted with a common one, the embedded watermark does not impact the output for the relevant query.
+For queries that partially retain trigger components from the original probe, such as Acronym-no_instruction and Spatial-imprecise, the watermark is retrieved at a limited rate.
+Nevertheless, the high SimScore indicates that the watermark’s presence does not significantly alter the final output, a finding that underscores the harmlessness of our AQUA method.
+Additional results are presented in Appendix C.2.
+5.5 Stealthiness of AQUA
+PCA Visualization.
+Inspired by prior work (Boler et al., 2022; Chen et al., 2024b; Yao et al., 2025), we observe that it is challenging to filter out watermarked images if their embeddings are highly similar to those of the original images in the dataset.
+Therefore, we randomly selected five watermarked images and their corresponding probe queries, along with 300 original images and normal queries, and visualized their embeddings using Principal Component Analysis (PCA) in Figure 5.
+These results indicate that AQUA maintains strong stealthiness while preserving high retrieval performance.
+VLM-based Filter as an Adaptive Adversary.
+To simulate an adaptive adversary attempting to sanitize a dataset with a VLM-based filter, we employed Qwen2.5-VL-Instruct (32B) to filter the dataset containing watermarked images.
+Results are averaged over 5 independent runs.
+In each trial, we randomly sampled 5,000 images from the MMQA dataset and randomly injected 10 different watermarked images.
+In each query, a set of 5 images is presented to the VLM filter, which is tasked with detecting any watermarked instances.
+The prompts used are provided in Appendix A.1.
+Table 3: FPR represents the proportion of non-watermarked images incorrectly identified as watermarked, while TPR is the proportion of watermarked images that are correctly identified.
+Methods Metrics Result
+$\textbf{AQUA}_{acronym}$ FPR $2.84\%$
+TPR $0\%$
+$\textbf{AQUA}_{spatial}$ FPR $2.36\%$
+TPR $0\%$
+Figure 5: PCA Visualization of Watermarked vs. Normal Images and Probe Queries vs. Normal Queries.
+The results in Table 5 show that AQUA’s watermarked images are highly resistant to the basic VLM-based filter, achieving 0% TPR while maintaining a low FPR for normal images.
+We further evaluate a more stringent VLM-based filter in Appendix C.3 (Table 12), where although the filter can successfully detect some watermarked images, it concurrently incurs a prohibitively high FPR on benign images.
+This confirms that AQUA’s images utilize semantic information within seemingly normal images as the verification signal, making any filtering approach face an inherent trade-off between detection rate and false positive rate.
+For visual comparisons confirming indistinguishability, please refer to Appendix A.3.
+Retrieval Ratio vs.
+Watermark Number.
+Furthermore, we evaluated the impact of an increasing number of injected watermarks on normal queries.
+Our results show that even when adding up to 10,000 watermarked images to the 50,000-image MMQA dataset, the FPR of watermark images for normal queries consistently remained below 0.1%.
+More results and figures can be found in Appendix C.3.
+5.6 Robustness of AQUA
+Image-Level Attacks.
+To evaluate the robustness of AQUA, we conducted experiments using the WAVES benchmark (An et al., 2024).
+For the experimental protocol, the attack ‘strength’ parameter is uniformly set to 1 across all watermark distortion and regeneration methods.
+A total of 50 watermarked images are selected for each technique, with the entire MMQA dataset serving as the original data corpus.
+All experimental results, generated by the Qwen2.5-VL-Instruct (7B) model, are presented in Table 4.
+The results indicate that images watermarked by AQUA sustain high retrieval rates and positive statistical verification outcomes following various image transformations, distortions, and regeneration attacks, which demonstrates the robustness of the proposed watermarking scheme.
+Table 4: Robustness under image attacks. “Regen_Both” denotes the sequential application of Regen_VAE and Regen_Diffusion.
+Attack $\mathbf{AQUA}_{acronym}$ $\mathbf{AQUA}_{spatial}$ Attack $\mathbf{AQUA}_{acronym}$ $\mathbf{AQUA}_{spatial}$
+[…]
+Rescale $1.026$ $99.33\%$ $1.355$ $95.78\%$ Regen_VAE $1.052$ $97.61\%$ $1.498$ $93.91\%$
+Rotate $1.071$ $98.54\%$ $1.613$ $89.80\%$ Regen_Diff. $1.036$ $98.17\%$ $1.502$ $94.33\%$
+Gaussian $1.068$ $99.00\%$ $1.459$ $91.21\%$ Regen_Both $1.037$ $96.55\%$ $1.516$ $87.39\%$
 […]
 […]
-InternVL3 Naive $2.86$ $27.11\%$ $0.41$ $4.56$ $17.12\%$ $0.65$
-Opt. $1.45$ $19.34\%$ $5.39e^{-3}$ $1.90$ $19.45\%$ $3.87e^{-3}$
-[…]
-$\textit{AQUA}_{spatial}$ $1.29$ $75.72\%$ $1.49e^{-50}$ $1.85$ $72.46\%$ $4.31e^{-26}$
-Qwen-VL -Chat Naive $2.86$ $15.79\%$ $0.59$ $4.56$ $5.71\%$ $0.91$
-Opt. $1.45$ $21.29\%$ $9.05e^{-3}$ $1.90$ $18.91\%$ $1.21e^{-3}$
-[…]
-[…]
-Qwen2.5- VL-Instruct Naive $2.86$ $38.15\%$ $0.25$ $4.56$ $15.87\%$ $0.86$
-Opt. $1.45$ $19.96\%$ $7.35e^{-3}$ $1.90$ $18.51\%$ $6.77e^{-3}$
-[…]
-$\textit{AQUA}_{spatial}$ $1.29$ $98.42\%$ $8.29e^{-72}$ $1.85$ $89.85\%$ $2.92e^{-49}$
-In this section, we evaluate the effectiveness of our AQUA.
-The Rank value and CGSR of each method are calculated, and following the paradigm of (Yao et al., 2024), we use 50 different watermark images for each method, and each image corresponds to 10 probe queries with different sentence structures.
-The experiment is repeated 10 times to obtain the p-values in the Table 1.
-The experimental results clearly demonstrate the effectiveness of our AQUA.
-Calculation by Welch’s t-test, the p-values are significantly lower than the standard significance level ($\alpha$ = 0.05).
-This allows us to confidently reject the null hypothesis $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, providing strong statistical evidence that the AQUA method enables reliable detection of the injected watermarks.
-Figure 4: TPR vs. FPR of two methods and two baselines.
-Figure 5: The relationship between p-value and query times.
-p-value vs.
-Query times.
-While all methods can eventually produce statistically significant results (i.e., low p-values) to reject the null hypothesis, the number of queries required to reach significance is a crucial factor in real-world applications—especially when each query may incur cost or be subject to limitations.
-To evaluate query efficiency, we measure how efficiently each method achieves statistical significance.
-Figure 5 shows that $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ can obtain a p-value less than the significance level within 30 queries, while the Opt.
-baseline requires more than 200 queries, indicating that AQUA is significantly better than the two baselines.
-FPR vs.
-TPR.
-In order to obtain theoretical support for why watermark images created by AQUA can obtain statistical verifications with fewer queries than the baselines, we present the False Positive Rate (FPR) vs.
-True Positive Rate (TPR) plot (Figure 5).
-FPRs are gained from measuring the generator’s (LLaVA-NeXT) output without watermark images in the database, and TPRs are obtained with 1, 2, 3, 5, and 10 watermark images for each probe query.
-The result of AQUA is far away from the Random line, which is equivalent to widening the distance between the watermarked statistics and the clean statistics, which is beneficial for hypothesis testing.
-Real-world Deployment.
-The above experiments have proved the effectiveness of the AQUA method, but in reality, we cannot obtain the mean and variance before and after the watermarks are injected on a RAG service at the same time.
-We can only get one mean and variance ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) from the suspected RAG service, so we propose a verification strategy with a predefined VSR’s reference distribution.
-We first characterize the reference distribution of a clean Multimodal RAG using mean and variance ($\mu_{clean}$, $\sigma^{2}_{clean}$), and the same with a watermarked one ($\mu_{wm}$, $\sigma^{2}_{wm}$).
-Subsequently, we can perform Welch’s t-test between ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) and two respective reference distributions.
-The null hypotheses ($\mathcal{H}_{0}$) for two hypothesis tests are: Suspect vs.
-Clean: $\mathcal{H}_{0}^{(1)}:\hat{\mu}_{suspect}<\mu_{clean}$ and Suspect vs.
-Watermarked: $\mathcal{H}_{0}^{(2)}:\hat{\mu}_{suspect}>\mu_{wm}$.
-To avoid a false accusation, the significance level $\alpha$ can be set to a very low value (e.g.
-$3e^{-5}$ in (Jovanović et al., 2025)).
-Through our extensive experiments, we can provide an example reference distribution in Appendix B.
-5.4 Harmlessness of AQUA
-Normal Query.
-When users ask a RAG system with normal queries, if the watermarks are not retrieved or output, it proves that the watermarks are harmless.
-We use normal queries (e.g.
-“What animals race in the Kentucky Derby?”) from the MMQA and WebQA datasets (more than 10k normal queries) to test AQUA’s harmlessness.
-Experiments show that when using a normal query within just one watermark image in the knowledge base, the retrieval rate of both $\textit{AQUA}_{acronym}$’s and $\textit{AQUA}_{spatial}$’s watermark images are 0%, and the CGSR is also 0% on four generators.
-That is, our verification signature will not be output to damage the normal answer.
-Table 2: Examples of relevant queries and corresponding results.
-Type Probe Query Relevant Query Rank SimScore $\uparrow$
-Acronym-
-replace
-[…]
-Acronym-
-no_instruction
-What is the subtitle of UGP? What is UGP? $1.07$ $70.18\%$
-Spatial-
-imprecise
-What fruit is the monkey holding like a phone? What is the monkey holding? $2.93$ $75.87\%$
-Relevant Query.
-Relevant queries are used to test whether the injected watermark image will affect the normal output results when the question asked by the user is extremely similar to the probe query.
-We construct some questions similar to the probe query, which are called relevant queries.
-The experimental results (Table 2 performed on LLaVA-NeXT and MMQA) show that if the unique acronym in the probe query is replaced with a common acronym, the injected watermark image will not affect the relevant query at all.
-Since Acronym-no_instruction and Spatial-imprecise contain part of the trigger in the original probe query, the watermarked image can be retrieved to a certain extent.
-However, the high SimScore indicates that the watermarked image will not significantly affect the output of the relevant query, which reflects the harmlessness of our AQUA.
-More results are shown in Appendix C.
-5.5 Stealthiness of AQUA
-(a) Image embeddings
-(b) Text embeddings
-Figure 6: PCA comparison on embeddings. Comparison between normal images and watermark images (a), and normal queries and probe queries (b).
-PCA Visualization.
-Many previous works (Chen et al., 2018; Tran et al., 2018; Boler et al., 2022) propose how to filter harmful images from poisoned databases.
-Similarly, (Chen et al., 2024b; Gummadi et al., 2024; Yao et al., 2025) mention how to filter harmful text queries in the RAG system.
-So, if the embeddings of watermark images and probe queries are similar to the embeddings of original images and normal queries in the dataset, detecting and filtering the watermark images is difficult for Adversary.
-We employ Principal Component Analysis (PCA) (Pearson, 1901; Hotelling, 1933) to visualize the embeddings of watermark images and their probe queries compared to those of original images and normal queries from the MMQA dataset (Fig.
-6).
-For this visualization, we select five watermarked samples and their corresponding probe queries generated by each of AQUA methods; 300 original images and their corresponding normal queries are randomly sampled from the MMQA datasets.
-These results indicate that AQUA maintains strong stealthiness while preserving high retrieval performance.
-Figure 7: The retrieval rate of watermarks under normal query as the number of injected watermark images increases.
-Retrieval Ratio vs.
-Watermark Number.
-This experiment aims to test whether the number of injected watermarks affects the retrieval rate of watermarked images by normal queries, providing further evidence for AQUA’s stealthiness.
-Figure 7 shows that as the number of injected watermark images increases, the retrieval ratio of watermark images constructed by our $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ are both less than 0.1% and does not increase significantly, further demonstrating the stealthiness of our approach.
-5.6 Robustness of AQUA
-{textblock*}
-0.63 (1pt,1pt)
-Table 3: The Rank and p-value values of the watermark image after the following transformations.
-Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
-Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
-Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
-Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
-Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
-Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
-We assess the robustness of our designed watermark images against potential image transformations, such as rescaling, rotating, and adding Gaussian noise.
-As in the setting of experiment 5.3, the p-value is obtained through 512 query trials.
-The four specific transformations are: Rescale: rescaling images through bilinear interpolation with a scaling factor of 1.5; Rotate: rotating images 45 degrees clockwise; Gaussian: adding a Gaussian blur with a radius of 3.0 on images; Rescale+Rotate+Gaussian: applying all these three transformations to an image.
-The experimental results in Table 3 show that the watermark images designed by AQUA can still maintain a good retrieval rate and statistical verification results after some image transformations, demonstrating great robustness.
-6 Discussion
-Limitation.
-Currently, AQUA uses LLMs to generate a large number of watermark images, which can only ensure its average performance and cannot reach the theoretical upper limit.
-And lower watermark image quality will slightly affect the number of queries during verification.
-Future Works.
-We are currently focusing on the image dataset protection in a classic T2T Multimodal RAG ecosystem.
-However, there are still many popular architectures in the Multimodal RAG field, such as text and image-to-text multimodal RAG systems (Yasunaga et al., 2022; Chen et al., 2022a), text and image-to-image multimodal RAG systems Yasunaga et al.
-(2022); Chen et al.
-(2022b); Shalev-Arkushin et al.
-(2025), and so on.
-Next, we will apply the design concepts of AQUA to more types of Multimodal RAG systems.
+System-Level Variations.
+In addition to image-level distortions and regeneration attacks, we further evaluate the stability of AQUA against practical deployment variations, such as benign system updates and corpus expansion.
+As detailed in Appendix C.4, altering the retrieval model (from CLIP to SigLIP (Zhai et al., 2023a)) or fine-tuning it yields only marginal changes to the retrieval Rank (Table 15).
+Furthermore, when the retrieval corpus is substantially enlarged by introducing 50k WebQA distractor images, the detection performance remains highly robust, achieving a CGSR of 96.76% for $\textbf{AQUA}_{acronym}$ and 91.85% for $\textbf{AQUA}_{spatial}$ (Table 16).
+These results confirm that AQUA maintains reliable performance under common real-world system modifications and data scaling.
+Generator-Level Adaptations.
+We further consider adversaries who attempt to circumvent the watermark at the generator level.
+First, we test whether frontier VLMs (Gemini-3-Pro, GPT-5.1-High, Qwen3-VL(Google DeepMind, 2025; OpenAI, 2025; Bai and others, 2025b) refuse to process watermarked images as “unnatural” inputs; no refusal behavior is observed across all models.
+Second, we simulate an adversarial fine-tuning attack using LoRA (Hu et al., 2022) to train the generator to specifically reject watermark-related queries.
+While CGSR decreases moderately, the watermark remains reliably detectable, and the fine-tuning incurs a significant utility degradation ($8.62\%$ on MMQA accuracy), making it an impractical strategy for adversaries.
+Full details are provided in Appendix C.4.1.
 6 Conclusion
-This research focuses on safeguarding image dataset copyright in T2T Multimodal RAG systems.
-We proposed AQUA, a watermarking framework that meets four design requirements: effectiveness, harmlessness, stealthiness, and robustness.
-Two complementary watermarking strategies in AQUA can protect the copyright of image datasets through statistical verification methods using only a few watermark images.
-Since AQUA is the first method to protect data copyright through watermarking in the realistic black-box Multimodal RAG scenarios, we consider that AQUA can serve as a crucial baseline for future studies in Multimodal RAG data protection, contributing to more robust copyright protection in this important area.
+This research focuses on safeguarding the copyright of image datasets in T2T Multimodal RAG systems.
+We propose AQUA, a watermarking framework that meets four design requirements: effectiveness, harmlessness, stealthiness, and robustness.
+The two complementary watermarking strategies in AQUA enable copyright protection for image datasets via statistical verification using only a few watermark images.
+Since AQUA is the first method to protect data copyright through watermarking in realistic black-box Multimodal RAG scenarios, AQUA can serve as a crucial baseline for future studies in Multimodal RAG data protection, contributing to more robust copyright protection in this important area.
+Acknowledgments
+We sincerely thank the anonymous reviewers for their detailed and constructive feedback.
+We are also grateful to the area chairs for their dedication in carefully handling the review process despite the significantly increased workload during ICLR 2026.
+This work is supported by Shanghai Engineering Research Center of Intelligent Vision and Imaging.
+We also thank members of the ASPIRE Lab at SIST, ShanghaiTech University, for helpful conversations during the course of this work.
 Appendix A Detailed Information of AQUA
 A.1 Prompt
-Acronym Prompt.
-Here is an example prompt (Figure 8) where we use LLM (Gemini-2.5-Pro) to build a large number of acronyms and their full names.
-{num_of_watermark} indicates the number of watermarks you want.
-Here is an example:
-(UGP, Unicorn Grammar Parser)
-Please create {num_of_watermark} pairs of uncommon acronyms and their
-full names based on this example.
-Figure 8: Prompt used to construct a large number of acronym pairs.
-SimScore Prompt.
-This is the prompt (Figure 9) where we ask LLM (Gemini-2.5-Pro) to determine whether two strings are semantically similar.
+Acronym Prompt Here is an example: (UGP, Unicorn Grammar Parser) Please create {num_of_watermark} pairs of uncommon acronyms and their full names based on this example.
+This is an example prompt where we use LLM (Gemini-2.5-Pro) to build a large number of acronyms and their full names. {num_of_watermark} indicates the number of watermarks you want.
+SimScore Prompt Determine the semantic similarity between the following two strings and give your score on a scale of 0-100: String 1: {Clean_Answer} String 2: {Watermark_Answer} Just answer with numbers.
+This is the prompt where we ask LLM (Gemini-2.5-Pro) to determine whether two strings are semantically similar.
 {Clean_Answer} and {Watermark_Answer} indicate the response of Multimodal RAG’s generator with the un-watermarked dataset and the watermarked one, respectively.
-Determine the semantic similarity between the following two strings and
-give your score on a scale of 0-100:
-String 1: {Clean_Answer}
-String 2: {Watermark_Answer}
-Just answer with numbers.
-Figure 9: Prompt used to obtain the SimScore between two strings.
+VLM-based Filter Prompt for $\textbf{AQUA}_{acronym}$ You are an OCR filter.
+As long as the picture contains unusual acronym, it is a watermark picture.
+Your task is to identify if any of these five images are suspected of being the watermarked images.
+If yes, please answer ’yes’ and indicate how many watermarked images there are.
+If no, please answer ’no’.
+VLM-based Filter Prompt for $\textbf{AQUA}_{spatial}$ You are a dataset watermark filter.
+Watermarked images are those that look unnatural or even defy common sense.
+You need to identify if any of these five images are suspected watermarks.
+If yes, please answer ’yes’ and indicate how many watermarked images there are.
+If no, please answer ’no’.
 A.2 Watermark Examples
-In this section, we will provide several classic watermark examples.
-The examples of $\textit{AQUA}_{acronym}$ method is in the Table 4 and $\textit{AQUA}_{spatial}$ in Table 6.
+The acronym images used by $\textbf{AQUA}_{acronym}$ can be generated through various means: they may be constructed manually, implemented programmatically via libraries, or produced with the assistance of Large Language Models (LLMs).
+There are no specific stylistic requirements for these images.
+They can range from simple black text on a white background, as exemplified above, to acronyms embedded within naturalistic images, as depicted in the Table 5.
 The sentence structure and wording of the probe query can be freely selected, but semantic consistency must be ensured.
-Table 4: Examples of watermark images designed by the $\textit{AQUA}_{acronym}$ method.
+Table 5: Examples of watermark images designed by the $\textbf{AQUA}_{acronym}$ method.
 Watermark Image Probe Query Verification Signature
@@
 Unicorn Grammar Parser
-1. Output the meaning of BLT.
-2. What does the acronym BLT stand for based on the images?
-3. Please output the subtitle of acronym BLT.
-Bouncing Llama Technologies
-1. Please state the full phrase that BLT abbreviates.
-2. Could you provide the expansion for the acronym BLT?
-3. What words do the letters in BLT represent?
-Quantum Walrus Xylophone
-1. What is the meaning of the acronym TPB?
-2. Could you spell out what TPB represents?
-3. Give me the full form of TPB.
-Temporal Platypus Bagpipe
+1. UGP is the name of a store. What is the full name of this store?
+2. Please output all the purple English letters on a black shop in a magic style
+Unicorn Grammar Parser
+1. Help me find a picture of a street with green plants. Please output the English words on the sign to the right of the green plants.
+2. Find an image with a green UGP acronym and output the English text on the street sign in full.
+Unicorn Grammar Parser
+1. Find a green book standing on the table and output the title of the book.
+2. What is the full name of a book titled UGP?
+Unicorn Grammar Parser
 1. What does XCO signify?
 2. Could you decode XCO?
-3. Give me the full form of TPB.
+3. Give me the full form of XCO.
 Xenon Cubist Ottoman
-Table 5: Examples of watermark images designed by the $\textit{AQUA}_{spatial}$ method.
+Here are three additional examples of the $\textbf{AQUA}_{spatial}$ method (Figure 7).
+Since the data provider adds the watermarks themselves, they have the flexibility to define watermarks that either slightly deviate from the dataset’s overall distribution or conform to it while containing subtle variations in detail.
+We provide a pipeline to guide data providers in achieving a balance between semantic distinctiveness and statistical naturalness for their specific datasets.
+The details of this pipeline are elaborated in Appendix A.4.
+Table 6: Examples of watermark images designed by the $\textbf{AQUA}_{spatial}$ method.
 Watermark Image Probe Query Verification Signature
@@
 AQUA
-Appendix B Reference Distribution
-$\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ need to use different means and variances to characterize their respective reference distributions.
-Since this reference distribution is related to the specific watermark image constructed and its performance, here we can give an example reference distribution through our extensive experiments:
-  • $\textit{AQUA}_{acronym}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.005, 0.02); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.6, 0.2)
-  • $\textit{AQUA}_{spatial}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.2, 0.2); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.55, 0.25)
-Appendix C More Results of Harmlessness of AQUA
+A.3 Visual Stealthiness Inspection
+To further demonstrate the high stealthiness and visual naturalness of AQUA, we provide an extended gallery of watermarked samples in Figure 6.
+The gallery is organized into three columns: the first column displays images generated via $\textbf{AQUA}_{acronym}$, the second column shows $\textbf{AQUA}_{spatial}$, and the third column presents original, benign images from the MMQA dataset for comparison.
+As illustrated, the watermarked images maintain high visual fidelity.
+To the naked eye, they appear indistinguishable from the standard benign images, verifying that the watermarks do not introduce noticeable artifacts or unnatural distortions that would compromise the visual quality of the dataset.
+AQUAacronym
+AQUAspatial
+MMQA (Normal)
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Figure 6: Extended Gallery of Watermarked vs.
+Normal Images.
+The figure presents 10 examples for each category, arranged in two columns per group.
+Columns 1-2: $\textbf{AQUA}_{acronym}$; Columns 3-4: $\textbf{AQUA}_{spatial}$; Columns 5-6: Original MMQA images.
+The high visual fidelity ensures indistinguishability from benign samples.
+A.4 Image Caption Template and Evaluation Pipeline
+We note that semantic distinctiveness in images is largely driven by object-level details, while naturalness is influenced by global factors such as composition, spatial arrangement, color, and background.
+Our image generation process is designed to preserve both aspects.
+To ensure semantic distinctiveness, we employ structured and controllable templates.
+We then populate these templates with concept pairs that exhibit low co-occurrence probabilities—these can be identified using methods such as word embedding similarity, BERT-based measures, or prompting large language models.
+Image Caption Template {number 1} {color 1} {object 1} {location/action/state} {number 2} {color 2} {object 2}
+Any part of this template can be used as the verification signal, for instance, “what {object 2} is”, or “the {number 1} of object 1”.
+If using the example from our paper, the template would be “{a} {red} {apple} {on the head of} {a} {} {dog}”, and the verification signal would be “apple”.
+To ensure statistical naturalness, we score the generated captions using pre-trained language models (e.g., BERT) and filter out those with high perplexity, which typically correspond to unnatural or ungrammatical sentences.
+Through this two-stage process, concept selection and language model-based filtering, the data provider can strike a balance between semantic distinctiveness and linguistic naturalness.
+For data providers who possess intimate knowledge of their datasets, intuitively constructed watermarks are often sufficient for effective copyright protection.
+The advanced pipeline serves as an optional extension to further enhance the performance of the AQUA method.
+We have omitted a detailed description of this pipeline from the main text to maintain simplicity and highlight the core effectiveness of our primary approach.
+Appendix B Real-world Deployment
+B.1 Reference Distribution Estimation
+Section 5.3 demonstrated the effectiveness of the AQUA method.
+However, in a real-world deployment, obtaining the simultaneous mean and variance of a RAG service before and after watermark injection is often infeasible.
+Typically, we can only observe a single set of statistics ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) from a suspected RAG service.
+To address this, we propose a verification strategy based on predefined reference distributions.
+We first characterize the reference distribution of a clean Multimodal RAG using its mean and variance ($\mu_{clean}$, $\sigma^{2}_{clean}$), and similarly for a watermarked system ($\mu_{wm}$, $\sigma^{2}_{wm}$).
+Subsequently, we perform Welch’s t-test to compare the observed statistics ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) against these two respective reference distributions.
+The null hypotheses ($\mathcal{H}_{0}$) for the two hypothesis tests are defined as follows:
+  • Suspect vs. Clean: $\mathcal{H}_{0}^{(1)}:\hat{\mu}_{suspect}<\mu_{clean}$
+  • Suspect vs. Watermarked: $\mathcal{H}_{0}^{(2)}:\hat{\mu}_{suspect}>\mu_{wm}$
+To mitigate the risk of false accusations, the significance level $\alpha$ is set to a stringent value (e.g., $3e^{-5}$ as in Jovanović et al. (2025)).
+Procedure for Establishing Reference Distributions.
+Since verification distributions vary across different datasets and watermark patterns, data providers should derive reference distributions specific to their own datasets.
+We outline the following general procedure:
+  1. 1.
+    Select a RAG System: Employ an arbitrary RAG system (either open-source or closed-source) as the testbed.
+  2. 2.
+    Clean Distribution: Use the non-watermarked dataset as the knowledge base.
+Following the watermark verification protocols in Sections 4.1 and 4.2, perform probe queries to obtain the VSR metrics, and calculate the mean ($\mu_{clean}$) and variance ($\sigma^{2}_{clean}$).
+  3. 3.
+    Watermarked Distribution: Replace the knowledge base with the watermarked dataset and repeat the same process to obtain ($\mu_{wm}$, $\sigma^{2}_{wm}$).
+Cross-System and Cross-Dataset Validation.
+To demonstrate the robustness of our verification mechanism across diverse settings, we conduct additional experiments using both MMQA and WebQA datasets with three different VLMs as the RAG backbone: LLaVA-NeXT, InternVL3, and Qwen2.5-VL-Instruct.
+As shown in Table 8, regardless of the specific RAG system or dataset employed, there remains a significant divergence between the clean and watermarked distributions, ensuring reliable verification across diverse environments.
+Table 8: Reference distribution statistics ($\mu_{clean}$, $\sigma^{2}_{clean}$, $\mu_{wm}$, $\sigma^{2}_{wm}$) across different RAG systems and datasets.
+A clear separation between clean and watermarked distributions is consistently observed.
+RAG System & Watermark Dataset $\mu_{clean}$ $\sigma^{2}_{clean}$ $\mu_{wm}$ $\sigma^{2}_{wm}$
+LLaVA-NeXT + $\textbf{AQUA}_{acronym}$ MMQA 0.01 0.02 0.60 0.20
+LLaVA-NeXT + $\textbf{AQUA}_{spatial}$ MMQA 0.20 0.20 0.55 0.25
+InternVL3 + $\textbf{AQUA}_{acronym}$ MMQA 0.01 0.01 0.86 0.13
+InternVL3 + $\textbf{AQUA}_{spatial}$ MMQA 0.12 0.18 0.75 0.21
+Qwen2.5-VL-Instruct + $\textbf{AQUA}_{acronym}$ MMQA 0.00 0.01 0.98 0.02
+Qwen2.5-VL-Instruct + $\textbf{AQUA}_{spatial}$ MMQA 0.07 0.13 0.97 0.09
+LLaVA-NeXT + $\textbf{AQUA}_{acronym}$ WebQA 0.01 0.02 0.77 0.03
+LLaVA-NeXT + $\textbf{AQUA}_{spatial}$ WebQA 0.04 0.11 0.86 0.15
+InternVL3 + $\textbf{AQUA}_{acronym}$ WebQA 0.01 0.01 0.79 0.19
+InternVL3 + $\textbf{AQUA}_{spatial}$ WebQA 0.12 0.18 0.76 0.23
+Qwen2.5-VL-Instruct + $\textbf{AQUA}_{acronym}$ WebQA 0.00 0.01 0.96 0.05
+Qwen2.5-VL-Instruct + $\textbf{AQUA}_{spatial}$ WebQA 0.07 0.13 0.89 0.10
+$\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ require distinct parameters to characterize their respective reference distributions, as these depend on the specific watermark construction and its triggering performance.
+Based on our extensive experiments, we provide the following conservative example reference distributions with intentionally broadened variances to accommodate potential system fluctuations in real-world deployments:
+  • $\textbf{AQUA}_{acronym}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = ($0.005$, $0.02$); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = ($0.6$, $0.2$)
+  • $\textbf{AQUA}_{spatial}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = ($0.2$, $0.2$); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = ($0.55$, $0.25$)
+As evident from Table 8, the actual separation between the clean and watermarked distributions is consistently distinct across all tested models and datasets, with the empirical statistics often showing even stronger separation than these conservative estimates.
+This confirms that once the watermark signal is successfully detected in a RAG system, the resulting VSR value typically diverges significantly from that of any non-infringing system, ensuring robust verification regardless of the specific RAG system or dataset configuration.
+B.2 Collision Avoidance in Multi-Data Provider Environments
+In a large-scale Retrieval-as-a-Service (RaaS) ecosystem involving multi-data providers, a critical concern is the potential for watermark collision, where triggers from different providers might conflict.
+AQUA addresses this challenge through a two-tiered approach: inherent probabilistic mitigation and a proposed centralized management architecture.
+Inherent Mitigation via Combinatorial Verification.
+As detailed in the methodology (Section 4), AQUA ensures uniqueness by requiring the injection and verification of a set of watermark images (e.g., $N=10$) coupled with a rigorous hypothesis testing framework.
+Unlike single-instance watermarking, our design relies on the joint probability of simultaneous retrieval and successful triggering across multiple distinct samples.
+Consequently, even without coordination, the probability of an accidental collision occurring concurrently across an entire set of watermark images from different providers is statistically negligible in real-world scenarios.
+Centralized Watermark Registry.
+For scenarios demanding absolute certainty—such as massive-scale commercial platforms—we propose the implementation of a Centralized Watermark Registry.
+Instead of relying solely on probabilistic independence, this approach positions the platform as a coordinating authority to explicitly manage watermark distribution.
+Functioning as a clearinghouse, the registry would require data providers to submit their intended triggers (e.g., acronyms or spatial configurations) prior to injection, allowing the system to cross-reference existing entries and preclude potential overlaps.
+To further streamline this process, the platform could proactively assign orthogonal trigger sets or exclusive namespaces to authorized providers.
+By transitioning from a probabilistic model to a deterministic registry, this architecture effectively eliminates accidental collisions, thereby simplifying dispute resolution and standardizing the verification protocol across the ecosystem.
+Appendix C More Experimental Results
+C.1 More Results of Effectiveness of AQUA
+Q: Why is Welch’s t-test the appropriate statistical method in this experimental setting?
+1) The standard Student’s t-test requires the assumption of equal variances (homogeneity of variance) between the two groups being compared.
+This assumption is not met in our analysis.
+For the datasets before and after watermarking, a given probe query retrieves different sets of images.
+Since the RAG generator’s output is conditioned on this retrieved context, the resulting VSR scores for the two groups are expected to have unequal variances.
+Therefore, Welch’s t-test, which does not assume equal variances, is the appropriate statistical method for our comparison.
+2) For each watermarked image, we conducted multiple detection trials using a set of similar yet distinct probe queries.
+This repeated experimentation ensures that the resulting data distribution meets the normality assumption required for Welch’s t-test.
+3) To ensure the stability and robustness of watermark detection in a practical deployment, we inject multiple watermarked images for a single probe query.
+For example, if the retriever returns the top-5 results, we can inject 10 watermarked images.
+This guarantees that for a given probe query, all images retrieved from the watermarked dataset are watermarked, while all images retrieved from the original dataset are normal.
+This design satisfies the independence assumption of Welch’s t-test.
+Two-proportion Z-test.
+While Welch’s t-test serves as a robust and powerful method for comparing the means of two independent groups, particularly when population variances are unequal, the two-proportion Z-test is an equally standard and widely applied statistical tool specifically tailored for comparing proportions.
+The rationale for employing the Z-test in our experimental setting is direct and compelling.
+The two-proportion Z-test is the canonical statistical method for evaluating whether an observed difference between two such proportions is statistically significant.
+Our experimental design, which involves two independent groups—the watermarked (experimental) and non-watermarked (control) datasets—and a large number of trials, perfectly aligns with the underlying assumptions of this test.
+It provides a rigorous framework for rejecting the null hypothesis that the performance is equivalent.
+Accordingly, we applied the two-proportion Z-test to our experimental data to quantitatively validate the efficacy of our watermarking scheme.
+The results of this analysis are presented below:
+Table 9: The table shows the p-values obtained from the Z-test.
+Models Methods MMQA WebQA
+LLaVA- NeXT Naive $0.45$ $0.91$
+Opt. $1.06e^{-3}$ $7.42e^{-2}$
+$\textbf{AQUA}_{acronym}$ $1.28e^{-273}$ $4.32e^{-173}$
+$\textbf{AQUA}_{spatial}$ $5.04e^{-43}$ $3.89e^{-38}$
+InternVL3 Naive $0.38$ $0.73$
+Opt. $4.97e^{-3}$ $6.19e^{-3}$
+$\textbf{AQUA}_{acronym}$ $2.89e^{-251}$ $3.91e^{-110}$
+$\textbf{AQUA}_{spatial}$ $4.31e^{-48}$ $7.73e^{-36}$
+Qwen-VL-Chat Naive $0.53$ $0.84$
+Opt. 5.19e35.19e{-3} $9.33e^{-2}$
+$\textbf{AQUA}_{acronym}$ $5.62e^{-127}$ $8.02e^{-86}$
+$\textbf{AQUA}_{spatial}$ $2.07e^{-52}$ $7.11e^{-27}$
+Qwen2.5-VL-Instruct Naive $0.32$ $0.81$
+Opt. $2.01e^{-2}$ $9.33e^{-3}$
+$\textbf{AQUA}_{acronym}$ $1.11e^{-175}$ $2.70e^{-133}$
+$\textbf{AQUA}_{spatial}$ $3.43e^{-71}$ $5.19e^{-47}$
+Effectiveness across Varying Parameter Scales.
+To comprehensively evaluate the scalability and generalizability of AQUA, we extended our experiments to include models with varying parameter scales, ranging from 2B to 38B.
+This analysis aims to verify the method’s applicability across diverse deployment scenarios, from resource-constrained edge devices to high-performance server-grade systems.
+We maintained the consistent experimental settings as detailed in Section 5.3.
+Given that the retrieval ranking is solely determined by the fixed retriever component, we focus our reporting on the $CGSR$, which reflects the generator’s ability to correctly decode the watermark.
+The results are presented in Table 10.
+Table 10: Effectiveness of AQUA across different model scales. We report the CGSRCGSR (%) on the MMQA dataset.
+Model $\textbf{AQUA}_{acronym}$ $\textbf{AQUA}_{spatial}$
+InternVL3-2B $70.53$ $63.76$
+Qwen3-VL-2B-Instruct $92.18$ $91.67$
+InternVL3.5-38B $90.71$ $84.33$
+Qwen3-VL-32B-Instruct $99.83$ $98.91$
+The empirical results demonstrate the robustness of AQUA regardless of model size:
+  • Low-resource Feasibility: The method remains effective even on lightweight models.
+Notably, the 2B-parameter Qwen3-VL achieves a CGSR exceeding $91\%$, outperforming several larger baselines.
+This indicates that AQUA is viable for deployment in low-resource environments.
+  • Scalability: As the model parameter count increases to the 30B+ range, the watermark triggering performance consistently improves, with Qwen3-VL-32B achieving near-perfect detection rates (up to $99.83\%$).
+This confirms that the watermarking mechanism aligns well with the enhanced reasoning capabilities of large-scale foundation models.
+C.2 More Results of Harmlessness of AQUA
 This section is a supplement to the experiment section on harmlessness of AQUA (Section 5.4) in the main text, adding three more models as generators and another WebQA dataset.
@@
 Rank SimScore $\uparrow$ Rank SimScore $\uparrow$
-LLaVA-NeXT Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
+LLaVA-NeXT Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
 Acronym-no_instruction $1.07$ $70.18\%$ $1.24$ $67.53\%$
 Spatial-imprecise $2.93$ $75.87\%$ $3.17$ $71.27\%$
-InternVL3 Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
+InternVL3 Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
 Acronym-no_instruction $1.07$ $71.28\%$ $1.2$ $68.29\%$
 Spatial-imprecise $2.93$ $68.92\%$ $3.17$ $63.31\%$
-Qwen-VL-Chat Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
+Qwen-VL-Chat Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
 Acronym-no_instruction $1.07$ $56.42\%$ $1.24$ $51.58\%$
 Spatial-imprecise $2.93$ $63.60\%$ $3.17$ $56.20\%$
-Qwen2.5-VL-Instruct Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
+Qwen2.5-VL-Instruct Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
 Acronym-no_instruction $1.07$ $82.85\%$ $1.24$ $78.51\%$
 Spatial-imprecise $2.93$ $78.23\%$ $3.17$ $69.82\%$
-Appendix D More Results of Robustness of AQUA
-In this section, we will test the robustness of our AQUA on more models in Table 8.
-Because robustness has little to do with the dataset, here we show the experimental results on the MMQA dataset as in the main text.
-Table 8: This table shows the Rank and p-value of watermark image after some common transformations.
-Models Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
-Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
-LLaVA-NeXT Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
-Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
-Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
-Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
-InternVL3 Rescale $1.03$ $4.85e^{-231}$ $1.36$ $4.85e^{-46}$
-Rotate $1.07$ $9.17e^{-112}$ $1.61$ $1.59e^{-31}$
-Gaussian $1.07$ $7.69e^{-229}$ $1.46$ $3.76e^{-39}$
-Rescale + Rotate + Gaussian $1.06$ $6.04e^{-216}$ $1.79$ $8.24e^{-27}$
-Qwen-VL-Chat Rescale $1.03$ $4.58e^{-154}$ $1.36$ $4.15e^{-45}$
-Rotate $1.07$ $8.08e^{-106}$ $1.61$ $6.82e^{-32}$
-Gaussian $1.07$ $5.26e^{-142}$ $1.46$ $9.14e^{-38}$
-Rescale + Rotate + Gaussian $1.06$ $5.12e^{-125}$ $1.79$ $6.78e^{-25}$
-Qwen2.5-VL-Instruct Rescale $1.03$ $3.19e^{-269}$ $1.36$ $7.68e^{-68}$
-Rotate $1.07$ $7.54e^{-238}$ $1.61$ $4.15e^{-51}$
-Gaussian $1.07$ $1.23e^{-251}$ $1.46$ $8.48e^{-62}$
-Rescale + Rotate + Gaussian $1.06$ $5.47e^{-268}$ $1.79$ $6.49e^{-46}$
+C.3 More Results of Stealthiness of AQUA
+Figure 7 illustrates the retrieval probability of watermarked images as a function of the number of injected images.
+Figure 7: The retrieval rate of watermarks under normal query as the number of injected watermark images increases.
+More Stringent VLM-based Filters Furthermore, we explored the use of other prompts to filter watermarked images from the AQUA method.
+To evaluate a more stringent adaptive attack scenario for the AQUA method, we simulated a stricter OCR-based filter.
+In contrast to the primary approach where a VLM assesses 5 images collectively, this filter evaluates images individually.
+An image is flagged if the VLM detects any presence of text within it.
+Intuitively, this method is expected to screen out the majority of $\textbf{AQUA}_{acronym}$’s text-based watermarks, but at the cost of an exceptionally high FPR on benign images.
+The prompt used and the experimental results are presented below:
+OCR Filter Simulator You are an image filter. You need to identify whether there is text in this picture. If yes, answer ’yes’, if not, answer ’no’.
+Similarly, we conducted an experiment to assess the VLM’s ability to discern plausibility from its internal knowledge.
+Single image is input to the model, which is prompted to determine whether the content is unnatural or semantically incongruous.
+The prompt and corresponding results are detailed below:
+Strange Image Filter from VLM’s Perspective You are an image filter. If you think this picture is unreasonable or unnatural, answer ’yes’, otherwise answer ’no’.
+Table 12: FPR represents the proportion of non-watermarked images incorrectly identified as watermarked, while TPR is the proportion of watermarked images that are correctly identified.
+Methods Metrics Exp. 1 Exp. 2 Exp. 3
+$\textbf{AQUA}_{acronym}$ FPR $67.60\%$ $61.44\%$ $72.02\%$
+TPR $\% 100100\%$ $100\%$
+$\textbf{AQUA}_{spatial}$ FPR $2.76\%$ $3.44\%$ $4.10\%$
+TPR $2\%$ $4\%$ $2\%$
+Our experimental results indicate that while a stringent VLM-based filter can remove a subset of the watermarked images, it concurrently incurs a prohibitively FPR on benign images.
+Given that an adaptive adversary’s primary objective is to augment their RAG system’s capabilities with the unauthorized database, adopting a filter that severely degrades the quality of legitimate data is an impractical strategy.
+This operational constraint for the adversary further underscores the stealthiness of our AQUA watermarking framework.
+C.4 More Results of Robustness of AQUA
+In this section, we provide a comprehensive evaluation of AQUA’s robustness.
+We examine two critical aspects: (1) resilience against adaptive adversaries who attempt to circumvent the watermark through model refusals or adversarial fine-tuning, and (2) stability under benign system updates, including component replacements and dataset drifts.
+C.4.1 Defense against Adversarial Adaptations
+We first address the concern of model-level adaptations by an adversary aiming to erase or evade the watermark triggers.
+Evaluation of Potential Refusals.
+To empirically address the concern that advanced VLMs might perceive watermarked images as “unnatural” and consequently refuse to generate responses, we conducted direct evaluations using state-of-the-art models, including Gemini-3-Pro, GPT-5.1-High, and Qwen3-VL.
+In our experiments, we observed no refusal behaviors.
+These models consistently processed the spatial queries as standard visual reasoning tasks, confirming that the watermark triggers remain compatible with high-level visual perception and do not trigger safety refusals.
+Robustness against Adversarial Fine-tuning.
+We simulated an adaptive adversary who attempts to bypass the watermark by fine-tuning the model to specifically refuse answering watermark-related queries.
+To construct the attack scenario, we fine-tuned the Qwen2.5-VL-7B-Instruct model on a mixed dataset designed to balance refusal behavior and general utility. Specifically, the training data included:
+  • Adversarial Samples: 100 newly generated task-specific adversarial samples paired with refusal responses (e.g., “I can’t help with that one because the watermark cues have to stay hidden.”).
+  • General Capabilities Data: 500 samples from COCO and 300 samples from Alpaca-GPT4 to maintain general visual reasoning capabilities.
+Configuration & Results. We employed LoRA targeting all linear layers combined with NEFTune noise injection (Jain et al., 2024) (hyperparameters listed in Table 13).
+Table 13: Hyperparameters for adversarial fine-tuning.
+Hyperparameter Value Hyperparameter Value
+Backbone Model
+Qwen2.5-VL-7B-Instruct
+LoRA Rank ($r$) 32
+LoRA Alpha ($\alpha$) 8 Target Modules All Linear
+NEFTune Noise $\alpha$ 5 Learning Rate $3e^{-5}$
+Batch Size 8 Epochs 1
+Warmup Ratio 0.03 Precision bf16
+We evaluated the fine-tuned model on watermark detectability (CGSR) and general utility (MMQA Accuracy on 2,000 normal QA pairs).
+As shown in Table 14, while CGSR saw a moderate decrease, the watermark remained effectively detectable.
+Crucially, this process caused a significant degradation in general utility ($8.62\%$).
+This indicates that forcing the model to reject watermarked images—which are distributionally similar to natural data—triggers catastrophic forgetting of general visual reasoning capabilities, acting as a strong deterrent to adversaries.
+Table 14: Results of adversarial fine-tuning. $\Delta$ denotes relative change.
+Metric Original Model Adversarially Fine-tuned Model $\Delta$
+$\textbf{AQUA}_{spatial}$ (CGSR) $98.42\%$ $89.34\%$ $9.23\%$
+$\textbf{AQUA}_{acronym}$ (CGSR) $99.61\%$ $87.18\%$ $12.48\%$
+Utility (MMQA Acc.) $27.25\%$ $24.90\%$ $8.62\%$
+C.4.2 Stability under System Updates and Drifts
+We empirically evaluate several common benign update scenarios (retriever replacement/fine-tuning and corpus expansion with distractors) and observe only marginal changes in Rank/CGSR, suggesting that AQUA remains stable under such deployment-time updates.
+Robustness of the Retriever Component.
+We evaluated the system’s performance under two distinct scenarios: architectural replacement and domain-specific adaptation.
+First, we replaced the default retriever with SigLIP (Zhai et al., 2023b) (siglip-so400m).
+Second, we fine-tuned the original retriever on the MMQA dataset.
+Crucially, to simulate a realistic deployment scenario where the model adapts to user queries, we utilized 1,000 image-question pairs for fine-tuning, rather than the conventional image-caption pairs.
+This setting better reflects the distribution shift encountered in real-world QA tasks.
+As shown in Table 15, the ranking metrics remain highly consistent, demonstrating that AQUA generalizes well across architectures and remains effective even after task-specific optimization.
+Table 15: Robustness results under retriever updates. We report the average RankRank (lower is better) for both methods.
+Retriever Setting Rank $\downarrow$
+$\textbf{AQUA}_{acronym}$ $\textbf{AQUA}_{spatial}$
+Standard (CLIP-ViT) $1.027$ $1.054$
+Replacement (SigLIP) $1.106$ $1.386$
+Fine-tuned (CLIP-ViT) $1.038$ $1.376$
+Resilience to Large-Scale Dataset Drifts.
+Finally, we tested the robustness of our approach against varying sizes and contents of the retrieval corpus.
+Based on the original MMQA dataset, we progressively added 10k to 50k images from the WebQA dataset as distractors.
+The results, detailed in Table 16, demonstrate that even with a significant increase in dataset size (nearly doubled), both the retrieval $Rank$ and $CGSR$ remain stable.
+This stability highlights the scalability of AQUA in handling dynamic, large-scale databases.
+Table 16: Performance stability under large-scale dataset drifts (adding WebQA distractors).
+Added Images +10k +20k +30k +40k +50k
+Retrieval Performance (Rank $\downarrow$)
+$\textbf{AQUA}_{acronym}$ $1.026$ $1.028$ $1.031$ $1.029$ $1.042$
+$\textbf{AQUA}_{spatial}$ $1.351$ $1.526$ $1.424$ $1.496$ $1.696$
+Semantic Decoding Performance (CGSR $\uparrow$)
+$\textbf{AQUA}_{acronym}$ $98.77\%$ $98.39\%$ $96.10\%$ $97.21\%$ $96.76\%$
+$\textbf{AQUA}_{spatial}$ $96.19\%$ $93.81\%$ $92.56\%$ $95.77\%$ $91.85\%$
+Appendix D Ethics Statement
+This work adheres to the ICLR Code of Ethics.
+In this study, no human subjects or animal experimentation were involved.
+All datasets used, including synthetic images, were sourced in compliance with relevant usage guidelines, ensuring no violation of privacy.
+We have taken care to avoid any biases or discriminatory outcomes in our research process.
+No personally identifiable information was used, and no experiments were conducted that could raise privacy or security concerns.
+We are committed to maintaining transparency and integrity throughout the research process.
+Appendix E Reproducibility Statement
+We have made every effort to ensure that the results presented in this paper are reproducible.
+All code has been uploaded as the supplemental materials to facilitate replication and verification.
+The experimental setup, including training steps, model configurations, is described in detail in the paper.
+Additionally, multimodal QA datasets, such as MMQA and WebQA, are publicly available, ensuring consistent and reproducible evaluation results.
+We believe these measures will enable other researchers to reproduce our work and further advance the field.
+Appendix F The Use of Large Language Models (LLMs)
+Large Language Models (LLMs) were used to aid in the writing and polishing of the manuscript.
+Specifically, we used an LLM to assist in refining the language, improving readability, and ensuring clarity in various sections of the paper.
+The model helped with tasks such as sentence rephrasing, grammar checking, and enhancing the overall flow of the text.
+It is important to note that the LLM was not involved in the ideation, research methodology, or experimental design.
+All research concepts, ideas, and analyses were developed and conducted by the authors.
+The contributions of the LLM were solely focused on improving the linguistic quality of the paper, with no involvement in the scientific content or data analysis.
+The authors take full responsibility for the content of the manuscript, including any text generated or polished by the LLM.
+We have ensured that the LLM-generated text adheres to ethical guidelines and does not contribute to plagiarism or scientific misconduct.